### v1

# MaxSup: Overcoming Representation Collapse in Label Smoothing

Abstract
Label Smoothing (LS) is widely adopted to curb overconfidence in neural network predictions and enhance generalization. However, previous research shows that LS can force feature representations into excessively tight clusters, eroding intra-class distinctions. More recent findings suggest that LS also induces overconfidence in misclassifications, yet the precise mechanism remained unclear. In this work, we decompose the loss term introduced by LS, revealing two key components: (i) a regularization term that functions only when the prediction is correct, and (ii) an error-enhancement term that emerges under misclassifications. This latter term compels the model to reinforce incorrect predictions with exaggerated certainty, further collapsing the feature space. To address these issues, we propose Max Suppression (MaxSup), which uniformly applies the intended regularization to both correct and incorrect predictions by penalizing the top-1 logit instead of the ground-truth logit. Through feature analyses, we show that MaxSup restores intra-class variation and sharpens inter-class boundaries. Extensive experiments on image classification and downstream tasks confirm that MaxSup is a more robust alternative to LS.  [^1] Code is available at: https://github.com/ZhouYuxuanYX/Maximum-Suppression-Regularization.
Machine Learning, ICML
Figure 1: Comparison of Label Smoothing (LS) and MaxSup. Left: MaxSup mitigates the intra-class compression induced by LS while preserving inter-class separability. Right: Grad-CAM visualizations show that MaxSup more effectively highlights class-discriminative regions than LS.
1 Introduction
Multi-class classification (Russakovsky et al., 2015; LeCun, 1998) conventionally relies on one-hot labels, implicitly treating each class as if it were completely orthogonal to every other. In reality, however, classes often share low-level attributes (Zeiler & Fergus, 2014; Silla & Freitas, 2011) or exhibit high-level semantic similarities (Chen et al., 2021; Yi et al., 2022; Novack et al., 2023), which makes the strict orthogonality assumption overly simplistic. This mismatch can lead to over-confident classifiers, ultimately reducing generalization (Guo et al., 2020).
To address overconfidence, Szegedy et al. (2016) introduced Label Smoothing (LS), blending a uniform distribution with the hard label to reduce the model’s certainty in the target class. LS has become a mainstay in both image recognition (He et al., 2016; Touvron et al., 2021; Liu et al., 2021; Zhou et al., 2022b) and neural machine translation (Gao et al., 2020; Alves et al., 2023), often improving accuracy and calibration (Müller et al., 2019). However, studies have also revealed that LS can produce overly tight clusters in the feature space (Kornblith et al., 2021; Sariyildiz et al., 2022; Xu & Liu, 2023), thereby lowering intra-class diversity and harming transferability (Feng et al., 2021). Meanwhile, Zhu et al. (2022) reported that LS inadvertently increases confidence in incorrect predictions, though the exact cause remained unclear.
In this paper, we show that LS’s training objective inherently contains an error-enhancement term that amplifies misclassified predictions, thus causing overconfident errors and tighter feature clusters (Section 3.1, Table 1). Extending Zhu et al. (2022), we define “overconfidence” in terms of the network’s top-1 prediction rather than calibration-based criteria. Our analysis further demonstrates that penalizing the ground-truth logit in misclassifications compresses the feature space, reducing intra-class variation (Table 2), as corroborated by Grad-CAM visualizations (Figure 3).
To overcome these limitations, we propose Max Suppression (MaxSup), which retains LS’s desirable regularization effect while eliminating its error-enhancement component. Rather than penalizing the ground-truth logit, MaxSup penalizes the largest logit, thus providing consistent regularization regardless of prediction correctness. By preventing ground-truth suppression during misclassification, MaxSup preserves richer intra-class variation and improves inter-class separability. As illustrated in Figure 1, this alleviates the compression and attentional shortcomings introduced by LS, leading to more robust feature representations. Extensive experiments on both image classification and semantic segmentation confirm that MaxSup not only alleviates intra-class collapse but also boosts final accuracy, enhances generalization, and strengthens transfer performance.
Our contributions are summarized as follows:
  • We present a logit-level analysis of Label Smoothing that unearths an ‘error-enhancement” term, revealing how LS inadvertently reinforces overconfidence in misclassified samples.
  • We propose Max Suppression (MaxSup), which preserves LS’s desired regularization while removing its detrimental error-enhancement component, thereby reducing intra-class compression and boosting both classification and downstream task performance.
2 Related Work
We survey regularization techniques before focusing on Label Smoothing (LS) and highlighting how MaxSup differs.
2.1 Regularization
Regularization enhances the generalization of deep neural networks by limiting model complexity through various strategies. Classical approaches such as $\ell_{2}$ (Krogh & Hertz, 1991) and $\ell_{1}$ (Zou & Hastie, 2005) regularization constrain large or sparse weights, respectively, while Dropout (Srivastava et al., 2014) randomly deactivates neurons to prevent feature co-adaptation. Among loss-based methods, Label Smoothing (LS) (Szegedy et al., 2016) redistributes probability mass away from the ground-truth class, improving both accuracy and calibration (Müller et al., 2019). Variants like Online Label Smoothing (OLS) (Zhang et al., 2021) and Zipf Label Smoothing (Zipf-LS) (Liang et al., 2022) adapt LS by considering the model’s evolving predictions, yet they still fail to address the fundamental issue that arises when the ground-truth logit is not maximal (Section 3.1, Table 1). Other loss-based regularizers, such as Confidence Penalty (Pereyra et al., 2017) and Logit Penalty (Dauphin & Cubuk, 2021), target different aspects of the output distribution. Confidence Penalty discourages overconfident predictions, whereas Logit Penalty minimizes the global $\ell_{2}$-norm of logits to improve feature separability (Kornblith et al., 2021). However, Logit Penalty can reduce intra-class variation, impairing transfer learning (Section 4).
Our MaxSup approach
MaxSup diverges from these methods by selectively penalizing only the top-1 logit $(z_{\textit{max}})$ rather than the ground-truth logit $(z_{gt})$. Unlike LS-based techniques, which can exacerbate errors by excessively shrinking $z_{gt}$ for misclassified samples, MaxSup uniformly applies regularization to all predictions, regardless of correctness. Consequently, it effectively sidesteps the error-enhancement issue, preserves richer intra-class diversity (Table 2), and sustains robust transfer performance across various datasets and architectures (Table 3).
2.2 Studies on Label Smoothing
Label Smoothing (LS) has also been extensively examined in the context of knowledge distillation. Yuan et al. (2020) showed that LS can act as a proxy for distillation, while Shen et al. (2021) explored its role within teacher–student frameworks. Chandrasegaran et al. (2022) further demonstrated that a low-temperature, LS-trained teacher can improve distillation performance. Meanwhile, Kornblith et al. (2021) found that LS tightens class clusters in feature space, reducing transfer performance. From a Neural Collapse (NC) perspective (Zhou et al., 2022a; Guo et al., 2024), LS drives the model toward rigid feature clusters, a phenomenon measured by Xu & Liu (2023) via a variability metric.
Comparison with existing LS techniques
Our primary objective is to mitigate the error-enhancement effect. Instead of refining a smoothed label, as in OLS or Zipf-LS, MaxSup directly penalizes the highest logit $z_{\textit{max}}$. This simple yet effective modification ensures uniform regularization even when $z_{gt}$ is not the top logit, thereby maintaining greater intra-class diversity and avoiding the performance degradation common to LS-based approaches (Section 3.2). Additionally, MaxSup integrates seamlessly with standard training pipelines, requiring no extra computational overhead beyond simply replacing LS.
3 Max Suppression Regularization (MaxSup)
We first partition the training objective into two components: the standard Cross-Entropy (CE) loss and a regularization term introduced by Label Smoothing (LS). By expressing LS in terms of logits (Theorem 3.3), we isolate two key factors: a regularization term that controls overconfidence and an error-enhancement term that enlarges the gap between the ground-truth logit $z_{gt}$ and any higher logits (Corollary 3.4, Equation 5), ultimately degrading performance. To address these shortcomings, we propose Max Suppression Regularization (MaxSup), which applies the penalty to the largest logit $z_{\textit{max}}$ rather than to $z_{gt}$ (Equation 8, Section 3.2). This shift delivers consistent regularization for both correct and incorrect predictions, preserves intra-class variation, and bolsters inter-class separability. Consequently, MaxSup mitigates the representation collapse found in LS, attains superior ImageNet-1K accuracy (Table 1), and improves transferability (Table 2, Table 3). The following sections elaborate on MaxSup’s formulation and its integration into the overall training pipeline.
3.1 Revisiting Label Smoothing
Label Smoothing (LS) is a regularization technique designed to reduce overconfidence by softening the target distribution. Rather than assigning probability $1$ to the ground-truth class and 00 to all others, LS redistributes a fraction $\alpha$ of the probability uniformly across all classes:
Definition 3.1.
For a classification task with $K$ classes, LS converts a one-hot label $\mathbf{y}\in\mathbb{R}^{K}$ into a soft label $\mathbf{s}\in\mathbb{R}^{K}$:
$s_{k}=(1-\alpha)y_{k}+\frac{\alpha}{K},$ (1)
where $y_{k}=\mathds{1}_{\{k=gt\}}$ denotes the ground-truth class. The smoothing factor $\alpha\in[0,1]$ reduces the confidence assigned to the ground-truth class and distributes $\tfrac{\alpha}{K}$ to other classes uniformly, thereby mitigating overfitting, enhancing robustness, and promoting better generalization.
To clarify the effect of LS on model training, we first decompose the Cross-Entropy (CE) loss into a standard CE term and an additional LS-induced regularization term:
Lemma 3.2.
Decomposition of Cross-Entropy Loss with Soft Labels.
$H(\mathbf{s},\mathbf{q})\;=\;H(\mathbf{y},\mathbf{q})\;+\;L_{\textit{LS}},$ (2)
where
$L_{\textit{LS}}\;=\;\alpha\,\biggl{(}H\Bigl{(}\tfrac{\mathbf{1}}{K},\mathbf{q}\Bigr{)}\;-\;H(\mathbf{y},\mathbf{q})\biggr{)}.$ (3)
Here, $\mathbf{q}$ is the predicted probability vector, $H(\cdot)$ denotes the Cross-Entropy, and $\frac{\mathbf{1}}{K}$ is the uniform distribution introduced by LS. This shows that LS adds a regularization term, $L_{\textit{LS}}$, which smooths the output distribution and helps to reduce overfitting. (See Appendix A for a formal proof.)
Building on Lemma 3.2, we next explicitly express $L_{\textit{LS}}$ at the logit level for further analysis.
Theorem 3.3.
Logit-Level Formulation of Label Smoothing Loss.
$L_{\textit{LS}}\;=\;\alpha\Bigl{(}z_{gt}\;-\;\frac{1}{K}\sum_{k=1}^{K}z_{k}\Bigr{)},$ (4)
where $z_{gt}$ is the logit corresponding to the ground-truth class, and $\tfrac{1}{K}\sum_{k=1}^{K}z_{k}$ is the average logit. Thus, LS penalizes the gap between $z_{gt}$ and the average logit, encouraging a more balanced output distribution and reducing overconfidence. (See Appendix B for the proof.)
The behavior of $L_{\textit{LS}}$ differs depending on whether $z_{gt}$ is already the maximum logit. Specifically, depending on whether the prediction is correct ($z_{gt}=z_{\textit{max}}$) or incorrect ($z_{gt}\neq z_{\textit{max}}$), we can decompose $L_{\textit{LS}}$ into two parts:
Corollary 3.4.
Decomposition of Label Smoothing Loss.
$L_{\textit{LS}}\;=\;\underbrace{\frac{\alpha}{K}\sum_{z_{m}<z_{gt}}\Bigl{(}z_{gt}-z_{m}\Bigr{)}}_{\text{Regularization}}\;+\;\underbrace{\frac{\alpha}{K}\sum_{z_{n}>z_{gt}}\Bigl{(}z_{gt}-z_{n}\Bigr{)}}_{\text{Error-Enhancement}},$ (5)
where $M$ and $N$ are the numbers of logits below and above $z_{gt}$, respectively ($M+N=K-1$). Note that the error-enhancement term vanishes when $z_{gt}=z_{\textit{max}}$.
  1. 1.
    Regularization: Penalizes the gap between $z_{gt}$ and any smaller logits, thereby moderating overconfidence.
  2. 2.
    Error-Enhancement: Penalizes the gap between $z_{gt}$ and larger logits, inadvertently increasing overconfidence in incorrect predictions.
Although LS aims to combat overfitting by reducing prediction confidence, its error-enhancement component can be detrimental for misclassified samples, as it widens the gap between the ground-truth logit $z_{gt}$ and the incorrect top logit. Concretely:
  1. 1.
    Correct Predictions $(z_{gt}=z_{\textit{max}})$: The error-enhancement term is zero, and the regularization term effectively reduces overconfidence by shrinking the gap between $z_{gt}$ and any smaller logits.
  2. 2.
    Incorrect Predictions $(z_{gt}\neq z_{\textit{max}})$: LS introduces two potential issues:
    • Error-Enhancement: Increases the gap between $z_{gt}$ and larger logits, reinforcing overconfidence in incorrect predictions.
    • Inconsistent Regularization: The regularization term lowers $z_{gt}$ yet does not penalize $z_{\textit{max}}$, which further impairs learning.
These issues with LS on misclassified samples have also been observed in prior work (Xia et al., 2024). By precisely identifying these two components (regularization vs. error-enhancement), we can design a more targeted solution.
Ablation Study on LS Components
To gauge the influence of each LS component, we conduct an ablation study on ImageNet-1K using a DeiT-Small model (Touvron et al., 2021) without Mixup or CutMix. As shown in Table 1, LS’s gains arise solely from the regularization term, whereas the error-enhancement term degrades performance. In contrast, our proposed Max Suppression Regularization (MaxSup) omits the error-enhancement effect and achieves higher accuracy by retaining the beneficial regularization.
Table 1: Ablation on LS components using DeiT-Small on ImageNet-1K (without CutMix or Mixup). “Regularization” denotes penalizing logits smaller than $z_{gt}$; “Error-Enhancement” penalizes logits larger than $z_{gt}$. MaxSup removes error-enhancement while retaining regularization.
Method Formulation Accuracy
Baseline 74.21
+ Label Smoothing $\tfrac{\alpha}{K}\sum_{z_{m}<z_{gt}}(z_{gt}-z_{m})$ 75.91
$\,+\,\tfrac{\alpha}{K}\sum_{z_{n}>z_{gt}}(z_{gt}-z_{n})$
+ Regularization $\tfrac{\alpha}{M}\sum_{z_{m}<z_{gt}}(z_{gt}-z_{m})$ 75.98
+ Error-Enhancement $\tfrac{\alpha}{N}\sum_{z_{n}>z_{gt}}(z_{gt}-z_{n})$ 73.63
+ Error-Enhancement $\alpha\,(z_{gt}-z_{\textit{max}})$ 73.69
+ MaxSup $\alpha\Bigl{(}z_{\textit{max}}-\tfrac{1}{K}\!\sum_{k=1}^{K}z_{k}\Bigr{)}$ 76.12
From Table 1, it is evident that LS’s overall accuracy boost is exclusively attributed to the regularization component, whereas error-enhancement consistently degrades performance (73.63% or 73.69%). Removing the error-enhancement term while keeping only the regularization improves accuracy slightly (75.98% vs. 75.91%). Finally, by avoiding error-enhancement entirely and preserving the helpful regularization, MaxSup achieves 76.12% accuracy—surpassing LS. This result underscores that MaxSup directly addresses LS’s primary shortcoming by consistently applying the intended regularization even when the model’s top-1 prediction is incorrect.
3.2 Max Suppression Regularization
Label Smoothing (LS) suffers from two main limitations: inconsistent regularization and error amplification. As discussed in Section 3.1 and illustrated in Table 1, LS penalizes the ground-truth logit $z_{gt}$ even for misclassified examples, thereby unnecessarily widening the gap between $z_{gt}$ and the incorrect top-1 logit. To address these critical shortcomings, we propose Max Suppression Regularization (MaxSup), which explicitly penalizes the largest logit $z_{\textit{max}}$ rather than $z_{gt}$. This crucial shift ensures uniform regularization across both correct and misclassified samples, effectively eliminating the error-amplification issue seen in LS (Table 1), and preserving the integrity of the ground-truth logit for more stable and robust learning performance.
Definition 3.5.
Max Suppression Regularization
We define the Cross-Entropy loss with MaxSup as follows:
$\underbrace{H(\mathbf{s},\mathbf{q})}_{\text{CE with Soft Labels}}\;=\;\underbrace{H(\mathbf{y},\mathbf{q})}_{\text{CE with Hard Labels}}\;+\;\underbrace{L_{\textit{MaxSup}}}_{\text{Max Suppression Loss}},$ (6)
where
$L_{\textit{MaxSup}}\;=\;\alpha\Bigl{(}H\bigl{(}\tfrac{\mathbf{1}}{K},\mathbf{q}\bigr{)}\;-\;H(\mathbf{y}^{\prime},\mathbf{q})\Bigr{)},$ (7)
and
[…]
so that $y^{\prime}_{k}=1$ identifies the model’s top-1 prediction and $y^{\prime}_{k}=0$ otherwise. Here, $H\bigl{(}\tfrac{\mathbf{1}}{K},\mathbf{q}\bigr{)}$ encourages a uniform output distribution to mitigate overconfidence, while $H(\mathbf{y}^{\prime},\mathbf{q})$ penalizes the current top-1 logit. By shifting the penalty from $z_{gt}$ (the ground-truth logit) to $z_{\textit{max}}$ (the highest logit), MaxSup avoids unduly suppressing $z_{gt}$ when the model misclassifies, thus overcoming Label Smoothing’s principal shortcoming.
Logit-Level Formulation of MaxSup
Building on the logit-level perspective introduced for LS in Section 3.1, we can express $L_{\textit{MaxSup}}$ as:
$L_{\textit{MaxSup}}\;=\;\alpha\Bigl{(}z_{\textit{max}}\;-\;\tfrac{1}{K}\sum_{k=1}^{K}z_{k}\Bigr{)},$ (8)
where $z_{\textit{max}}=\max_{k}\{z_{k}\}$ is the largest (top-1) logit, and $\tfrac{1}{K}\sum_{k=1}^{K}z_{k}$ is the mean logit. Unlike LS, which penalizes the ground-truth logit $z_{gt}$ and may worsen errors in misclassified samples, MaxSup shifts the highest logit uniformly, thus providing consistent regularization for both correct and incorrect predictions. As shown in Table 1, this approach eliminates LS’s error-amplification issue while preserving the intended overconfidence suppression.
Comparison with Label Smoothing
MaxSup fundamentally differs from LS in handling correct and incorrect predictions. When $z_{gt}=z_{\textit{max}}$, both LS and MaxSup similarly reduce overconfidence. However, when $z_{gt}\neq z_{\textit{max}}$, LS continues to shrink $z_{gt}$, widening the gap with the incorrect logit, whereas MaxSup penalizes $z_{\textit{max}}$, preserving $z_{gt}$ from undue suppression. As illustrated in Figure 3, this allows the model to recover from mistakes more effectively and avoid reinforcing incorrect predictions.
Gradient Analysis
To understand MaxSup’s optimization dynamics, we compute its gradients with respect to each logit $z_{k}$. Specifically,
$\frac{\partial L_{\textit{MaxSup}}}{\partial z_{k}}\;=\;\begin{cases}\alpha\Bigl{(}1-\tfrac{1}{K}\Bigr{)},&\text{if }k=\arg\max(\mathbf{q}),\\ -\tfrac{\alpha}{K},&\text{otherwise}.\end{cases}$ (9)
Thus, the top-1 logit $z_{\textit{max}}$ is reduced by $\alpha\bigl{(}1-\tfrac{1}{K}\bigr{)}$, while all other logits increase slightly by $\tfrac{\alpha}{K}$. In misclassified cases, the ground-truth logit $z_{gt}$ is therefore spared from penalization, thereby avoiding the error-amplification issue seen in LS. For completeness, Appendix A provides a full derivation of these gradients, and Figure 2 compares the resulting logit distributions under different regularizers.
Behavior Across Different Samples
MaxSup applies a dynamic penalty that depends on the model’s current predictions. For high-confidence, correctly classified examples, it behaves similarly to LS by reducing overconfidence, thus effectively mitigating overfitting. In contrast, for misclassified or uncertain samples, MaxSup specifically suppresses the incorrect top-1 logit, further safeguarding the ground-truth logit $z_{gt}$. This selective strategy preserves an accurate representation of the true class while actively discouraging the propagation of errors. As shown in Section 5.1 and Table 5, this promotes more robust decision boundaries and ultimately leads to stronger generalization performance.
Theoretical Insights and Practical Benefits
MaxSup provides both theoretical and practical advantages compared to LS. Whereas LS applies a uniform penalty to the ground-truth logit regardless of correctness, MaxSup focuses on penalizing only the most confident logit $z_{\textit{max}}$. This dynamic adjustment prevents error accumulation in misclassifications, thereby ensuring more stable convergence. As a result, MaxSup achieves stronger generalization, exhibits greater robustness to label noise, and performs well on challenging datasets. Moreover, as shown in Section 4, MaxSup preserves higher intra-class diversity, which substantially improves transfer learning performance (Table 3) and yields more interpretable activation maps (Figure 3).
4 Analysis of MaxSup’s Learning Benefits
MaxSup simultaneously promotes inter-class separability and intra-class variation, both essential for robust classification and effective feature transfer. In this section, we explore how MaxSup achieves these objectives and contrast its effectiveness with alternative regularization methods.
4.1 Intra-Class Variation and Transferability
As noted in Section 3.1, Label Smoothing (LS) primarily restricts overconfidence when the ground-truth class is correctly predicted, inadvertently causing error enhancement for misclassified samples. This selective penalty can overly compress intra-class diversity. In contrast, MaxSup uniformly penalizes the top-1 logit in both correct and incorrect cases, eliminating LS’s error-enhancement component and thus preserving more fine-grained distinctions within each class. Table 2 compares intra-class variation $\bar{d}_{\text{within}}$ and inter-class separability $R^{2}$ (Kornblith et al., 2021) for a ResNet-50 model trained on ImageNet-1K. Although all regularization strategies reduce $\bar{d}_{\text{within}}$ relative to the baseline, MaxSup shows the smallest reduction, implying stronger retention of within-class variability—often correlated with improved generalization and transferability. The benefits of this richer intra-class structure appear clearly in Table 3, where linear transfer performance on CIFAR-10 is reported. Although LS and Logit Penalty improve ImageNet accuracy, they diminish transfer accuracy by over-suppressing informative features. In contrast, MaxSup preserves transfer performance near that of the baseline, suggesting it retains crucial, discriminative features that generalize effectively to downstream tasks.
Table 2: Metrics for feature representation quality using ResNet-50 trained on ImageNet-1K. We report intra-class variation ($\bar{d}_{\text{within}}$) and inter-class separability ($R^{2}$), both of which benefit from higher values. Although all methods reduce $\bar{d}_{\text{within}}$ relative to the baseline, MaxSup preserves the most within-class diversity.
Method $\bar{d}_{\text{within}}\uparrow$ $R^{2}\uparrow$
Train Val Train Val
Baseline 0.3114 0.3313 0.4025 0.4451
LS 0.2632 0.2543 0.4690 0.4611
OLS 0.2707 0.2820 0.5943 0.5708
Zipf’s 0.2611 0.2932 0.5522 0.4790
MaxSup 0.2926 0.2998 0.5188 0.4972
Logit Penalty 0.2840 0.3144 0.6448 0.6024
4.2 Impact of Logit Regularization
Different regularization methods impose distinct constraints on the logit space, thereby shaping the model’s representational capacity (Kornblith et al., 2021). Among these approaches, Logit Penalty and MaxSup both act directly on logits but differ fundamentally in how they apply regularization. Logit Penalty operates by minimizing the $\ell_{2}$-norm of the entire logit vector, causing a global reduction in logit magnitudes that often induces sparsity. This uniform shrinkage can limit intra-class variation, thereby weakening the model’s ability to transfer features to downstream tasks. In contrast, MaxSup targets only the largest (top-1) logit, nudging it closer to the average logit. By selectively penalizing only the most confident prediction, MaxSup avoids universal shrinkage and preserves richer intra-class diversity, a property crucial for effective transferability. Figure 2 illustrates the distribution of logits under various regularizers. Logit Penalty yields a narrower logit range, reflecting excessive sparsity and aligning with its lower transfer performance (Table 3). By comparison, MaxSup maintains broader logit distributions, thereby retaining the fine-grained feature distinctions needed to excel on downstream tasks.
Table 3: Validation performance on CIFAR-10 with a linear probe using $l_{2}$-regularized multinomial logistic regression. Although Label Smoothing and Logit Penalty improve ImageNet accuracy, they substantially degrade transfer accuracy compared to MaxSup.
Method Linear Transfer Acc.
Baseline 0.8143
Label Smoothing 0.7458
Logit Penalty (Dauphin & Cubuk, 2021) 0.7242
MaxSup 0.8102
Figure 2: Logit density plots under three different regularization strategies: MaxSup, Logit Penalty, and standard Cross Entropy. Logit Penalty induces a narrower logit distribution, reflecting excessive shrinkage that reduces intra-class variation. By contrast, MaxSup preserves a broader range of logits and thus richer representations.
5 Experiments
5.1 Evaluation on ImageNet Classification
In this section, we assess the effectiveness of MaxSup on ImageNet-1K, comparing its performance against standard Label Smoothing and related variants.
5.1.1 Experiment Setup
Model Training Configurations
We conduct extensive experiments with both CNN and Transformer models, including the ResNet family (He et al., 2016), MobileNetV2 (Sandler et al., 2018), and DeiT-Small (Touvron et al., 2021), all thoroughly evaluated on the large-scale ImageNet dataset (Krizhevsky et al., 2012) for comprehensive performance analysis.
For ResNet Series models, we train for 200 epochs using stochastic gradient descent (SGD) with momentum 0.9, a weight decay of $1\times 10^{-4}$, and a batch size of 2048. The initial learning rate is set to 0.85 and scheduled via cosine annealing. [^2] Additional training hyperparameters follow the FFCV training scripts in https://github.com/libffcv/ffcv. See Appendix E for further details on training setups. We also evaluate ResNet-based CNNs on CIFAR-100. Here, we use an initial learning rate of 0.1, reducing it by a factor of 5 at the 60th, 120th, and 160th epochs. We train for 200 epochs with a batch size of 128, weight decay of $5\times 10^{-4}$, and Nesterov momentum set to 0.9. For DeiT-Small, we employ the official implementation and train from scratch without knowledge distillation. Although the original DeiT paper emphasizes distillation, we exclude it to provide a clearer, unbiased assessment of MaxSup’s contributions. We also omit CutMix and Mixup to retain the same optimization objective.
Hyperparameters for Compared Methods
We compare Max Suppression Regularization against multiple Label Smoothing variants, including Zipf Label Smoothing (Liang et al., 2022) and Online Label Smoothing (Zhang et al., 2021). When official implementations are available, we use them directly; otherwise, we follow the respective papers’ descriptions closely to ensure fair comparisons. All training hyperparameters are kept identical to those of the baseline models, except for algorithm-specific settings and necessary adjustments. Additionally, we employ a linearly increasing $\alpha$ scheduler, which generally benefits training and stability; see Appendix F for details. This scheduler is applied to both MaxSup and standard Label Smoothing by default to maintain consistency.
5.1.2 Experiment Results
ConvNet Comparison
Table 4 summarizes the performance of MaxSup alongside various smoothing and self-distillation methods on both ImageNet and CIFAR-100. Across all tested convolutional architectures, MaxSup achieves the highest accuracy among label smoothing–based regularizers. By contrast, OLS (Zhang et al., 2021) and Zipf-LS (Liang et al., 2022) yield less consistent gains, suggesting their reported empirical benefits may depend heavily on specific training schedules. In our reproductions of OLS and Zipf-LS, we follow the authors’ original codebases and method-specific hyperparameters but do not adopt their complete training recipes. For example, the OLS paper uses a step learning-rate scheduler over 250 epochs with an initial rate of 0.1, while Zipf-LS trains for 100 epochs under a separate set of hyperparameters. Our results underscore the robustness of MaxSup across different training setups, in contrast to the more scheme-dependent improvements noted for OLS and Zipf-LS.
Table 4: Comparison of classic convolutional neural networks on ImageNet and CIFAR-100. Results are reported as “mean ± std” (percentage). Bold entries highlight the best performance; underlined entries mark the second best. (Methods with denote code adaptations from official repositories; see text for details.)
Method ImageNet CIFAR-100
ResNet-18 ResNet-50 ResNet-101 MobileNetV2 ResNet-18 ResNet-50 ResNet-101 MobileNetV2
Baseline 69.09±0.12 76.41±0.10 75.96±0.18 71.40±0.12 76.16±0.18 78.69±0.16 79.11±0.21 68.06±0.06
Label Smoothing 69.54±0.15 76.91±0.11 77.37±0.15 71.61±0.09 77.05±0.17 78.88±0.13 79.19±0.25 69.65±0.08
[…]
$\text{OLS}^{*}$ 69.45±0.15 77.23±0.21 77.71±0.17 71.63±0.11 77.33±0.15 78.79±0.12 79.25±0.15 68.91±0.11
MaxSup 69.96±0.13 77.69±0.07 78.18±0.12 72.08±0.17 77.82±0.15 79.15±0.13 79.41±0.19 69.88±0.07
Logit Penalty 68.48±0.10 76.73±0.10 77.20±0.15 71.13±0.10 76.41±0.15 78.90±0.16 78.89±0.21 69.46±0.08
DeiT Comparison
Table 5 compares various regularization techniques for DeiT-Small on ImageNet. MaxSup achieves an accuracy of $76.49\%$, surpassing Label Smoothing by $0.41$ percentage points. Label Smoothing variants such as Zipf’s and OLS yield only marginally higher or comparable performance relative to standard LS, suggesting that these approaches may be less effective for vision transformer architectures—potentially due to their reliance on extensive data augmentation schemes. In contrast, MaxSup consistently outperforms both standard LS and its variants, indicating its stronger ability to enhance feature representations without additional data manipulations. These findings underscore MaxSup’s robustness across distinct model architectures, especially in settings where other regularization methods show limited effectiveness.
Table 5: Accuracy (%) comparison on DeiT-Small (Touvron et al., 2021) using different Label Smoothing variants. Results are reported as “mean ± std”; parentheses indicate absolute improvement over the baseline.
Method Accuracy (Mean) Std
Baseline 74.39 0.19
Label Smoothing 76.08 ($+1.69$) 0.16
Zipf-LS 75.89 ($+1.50$) 0.26
OLS 76.16 ($+1.77$) 0.18
MaxSup 76.49 ($+2.10$) 0.12
Table 6: Semantic segmentation results on the ADE20K validation set. Models are pretrained on ImageNet-1K and then fine-tuned with UperNet (Xiao et al., 2018) for enhanced performance. We report mean Intersection over Union (mIoU) under multi-scale (MS) testing for comprehensive evaluation.
Backbone Method mIoU (MS)
DeiT-Small (Touvron et al., 2021) Baseline 42.1
Label Smoothing 42.4 ($+0.3$)
MaxSup 42.8 ($+0.7$)
Figure 3: Class activation maps generated by Grad-CAM (Selvaraju et al., 2019) for DeiT-Small models trained with MaxSup (2nd row), Label Smoothing (3rd row), and a standard Baseline (4th row). The first row shows the original images. Compared with Label Smoothing, MaxSup more effectively suppresses distractions from non-target objects and preserves key features of the target class, thereby reducing instances in which the model partially or completely focuses on irrelevant regions.
5.2 Evaluation on Semantic Segmentation
To further assess the transferability of MaxSup to downstream tasks, we evaluate its performance on semantic segmentation using the MMSegmentation framework. [^3] https://github.com/open-mmlab/mmsegmentation Specifically, we employ the UperNet architecture (Xiao et al., 2018) with a DeiT-Small backbone, trained on ADE20K. We compare backbones trained with Label Smoothing and MaxSup (on ImageNet-1K) against a baseline, following the same setup as in Section 5.1.2. During fine-tuning, all models use a standard cross-entropy loss. As shown in Table 6, MaxSup achieves a mean Intersection over Union (mIoU) of $42.8\%$, surpassing the $42.4\%$ obtained with Label Smoothing. These findings further highlight the improved feature representations afforded by MaxSup in downstream tasks such as semantic segmentation.
5.3 Visualization via Class Activation Maps
To assess how MaxSup influences model decision-making relative to Label Smoothing (LS), we employ Gradient-weighted Class Activation Mapping (Grad-CAM) (Selvaraju et al., 2019). Grad-CAM produces class-discriminative localization maps that highlight the regions most relevant to each classification decision.
We conduct comprehensive experiments on the DeiT-Small model under three distinct training setups: MaxSup (second row), Label Smoothing (third row), and standard Cross-Entropy (CE) as a baseline (fourth row). As illustrated in Figure 3, MaxSup-trained models exhibit a distinct advantage in effectively mitigating distractions caused by non-target salient objects in the background (e.g., a pole in the ‘Bird’ image, a tube in the ‘Goldfish’ image, and a cap in the ‘House Finch’ image). By contrast, LS-trained models often lose focus or erroneously attend to background objects, further reflecting the detrimental influence of LS’s Error-Enhancement term and its impact on feature learning (Please see Figures 3).
Moreover, MaxSup preserves a wider range of relevant object features, as evidenced in the ‘Shark’ and ‘Monkey’ examples, where LS-trained models fail to capture key details (fins and tails). These observations align with the analysis in Appendix G, underscoring that MaxSup better retains rich intra-class information. Consequently, MaxSup-trained models yield more accurate and reliable classifications by leveraging these detailed feature representations.
5.4 Quantitative Analysis of CAM Overlays
To further quantify the benefits of MaxSup, we measure precision and recall for Grad-CAM overlays on target object regions. Our results indicate that MaxSup consistently attains higher localization precision while maintaining or improving recall, leading to more accurate and comprehensive feature activation. Qualitative assessments similarly reveal that MaxSup-trained models exhibit both a more focused and more extensive attention map, resulting in more reliable and interpretable predictions. These findings demonstrate that MaxSup not only enhances feature representation but also contributes to more trustworthy AI systems.
6 Conclusion
In this work, we carefully examined the root causes of Label Smoothing’s (LS) shortcomings and introduced Max Suppression Regularization (MaxSup) as a targeted remedy. Our analysis shows that LS can unintentionally promote overconfidence in misclassified samples by applying insufficient regularization to erroneous top-1 logits. In contrast, MaxSup effectively addresses this by consistently penalizing the most confident logit, regardless of prediction correctness. Through extensive experiments and in-depth analyses, we demonstrate that MaxSup not only improves accuracy but also preserves richer intra-class variation and significantly enhances inter-class separability. Consequently, models trained with MaxSup capture finer-grained information about individual samples, ultimately leading to stronger transfer learning capabilities. Class activation maps further reveal that MaxSup directs model attention more accurately toward salient parts of target objects, effectively mitigating distractions from background elements.
Limitations & Future Work
Although our findings validate MaxSup’s effectiveness, several directions merit further investigation. Prior work (Müller et al., 2019) shows that teachers trained with LS can degrade performance in Knowledge Distillation (Hinton, 2015), and Guo et al. (2024) suggests LS accelerates convergence via conditioning number analysis. Future research could explore MaxSup’s impact on Knowledge Distillation workflows and its influence on training convergence. Additionally, recent studies (Súkeník et al., 2024; Garrod & Keating, 2024) indicate that $\ell_{2}$ regularization biases final layer features and weights toward lower-rank solutions than those typically associated with neural collapse. Investigating how MaxSup interacts with these low-rank biases and whether it leads to similarly optimal or novel solutions is another intriguing avenue for future work.
Impact Statement
This paper advances machine learning regularization techniques through an in-depth analysis of Label Smoothing and a proposed alternative, Max Suppression Regularization. By reinforcing generalization and transferability on widely used computer vision tasks, MaxSup has the potential to benefit a broad array of real-world applications, improving robustness without introducing known negative consequences. Thus, we anticipate a net positive impact, offering practitioners a clearer understanding of network regularization and access to more effective training methodologies.
Acknowledgements
This research was supported by the University of Washington Graduate School’s GSFEI Top Scholar Award and the Start-Up Funding awarded to Prof. Zhi-Qi Cheng. Additional support was provided by ELSA – European Lighthouse on Secure and Safe AI (No. 101070617).
Appendix A Proof of Lemma 3.2
Proof.
We aim to demonstrate the validity of Lemma 3.2, which states:
$H(\mathbf{s},\mathbf{q})=H(\mathbf{y},\mathbf{q})+L_{LS}$ (10)
where $L_{LS}=\alpha\left(H\left(\frac{\mathbf{1}}{K},\mathbf{q}\right)-H(\mathbf{y},\mathbf{q})\right)$
Let us proceed with the proof:
We begin by expressing the cross-entropy $H(\mathbf{s},\mathbf{q})$:
$H(\mathbf{s},\mathbf{q})=-\sum_{k=1}^{K}s_{k}\log q_{k}$ (11)
In the context of label smoothing, $s_{k}$ is defined as:
$s_{k}=(1-\alpha)y_{k}+\frac{\alpha}{K}$ (12)
where $\alpha$ is the smoothing parameter, $y_{k}$ is the original label, and $K$ is the number of classes.
Substituting this expression for $s_{k}$ into the cross-entropy formula:
$H(\mathbf{s},\mathbf{q})=-\sum_{k=1}^{K}\left((1-\alpha)y_{k}+\frac{\alpha}{K}\right)\log q_{k}$ (13)
Expanding the sum:
$H(\mathbf{s},\mathbf{q})=-(1-\alpha)\sum_{k=1}^{K}y_{k}\log q_{k}-\frac{\alpha}{K}\sum_{k=1}^{K}\log q_{k}$ (14)
We recognize that the first term is equivalent to $(1-\alpha)H(\mathbf{y},\mathbf{q})$, and the second term to $\alpha H(\frac{\mathbf{1}}{K},\mathbf{q})$. Thus:
$H(\mathbf{s},\mathbf{q})=(1-\alpha)H(\mathbf{y},\mathbf{q})+\alpha H\left(\frac{\mathbf{1}}{K},\mathbf{q}\right)$ (15)
Rearranging the terms:
$H(\mathbf{s},\mathbf{q})=H(\mathbf{y},\mathbf{q})+\alpha\left(H\left(\frac{\mathbf{1}}{K},\mathbf{q}\right)-H(\mathbf{y},\mathbf{q})\right)$ (16)
We can now identify $H(\mathbf{y},\mathbf{q})$ as the original cross-entropy loss and $L_{LS}=\alpha\left(H\left(\frac{\mathbf{1}}{K},\mathbf{q}\right)-H(\mathbf{y},\mathbf{q})\right)$ as the Label Smoothing loss.
Therefore, we have demonstrated that:
$H(\mathbf{s},\mathbf{q})=H(\mathbf{y},\mathbf{q})+L_{LS}$ (17)
with $L_{LS}$ as defined in the lemma. It is noteworthy that the original cross-entropy loss $H(\mathbf{y},\mathbf{q})$ remains unweighted by $\alpha$ in this decomposition, which is consistent with the statement in Lemma 3.2 ∎
Appendix B Proof of Theorem 3.3
Proof.
We aim to prove the equation:
$L_{\textit{LS}}=\alpha(z_{gt}-\frac{1}{K}\sum^{K}_{k=1}z_{k})$ (18)
Let $\mathbf{s}$ be the smoothed label vector and $\mathbf{q}$ be the predicted probability vector. We start with the cross-entropy between $\mathbf{s}$ and $\mathbf{q}$:
$H(\mathbf{s},\mathbf{q})=-\sum_{k=1}^{K}s_{k}\log q_{k}$ (19)
With label smoothing, $s_{k}=(1-\alpha)y_{k}+\frac{\alpha}{K}$, where $\mathbf{y}$ is the one-hot ground truth vector and $\alpha$ is the smoothing parameter. Substituting this:
$H(\mathbf{s},\mathbf{q})=-\sum_{k=1}^{K}[(1-\alpha)y_{k}+\frac{\alpha}{K}]\log q_{k}$ (20)
Expanding:
$H(\mathbf{s},\mathbf{q})=-(1-\alpha)\sum_{k=1}^{K}y_{k}\log q_{k}-\frac{\alpha}{K}\sum_{k=1}^{K}\log q_{k}$ (21)
Since $\mathbf{y}$ is a one-hot vector, $\sum_{k=1}^{K}y_{k}\log q_{k}=\log q_{gt}$, where $gt$ is the index of the ground truth class:
$H(\mathbf{s},\mathbf{q})=-(1-\alpha)\log q_{gt}-\frac{\alpha}{K}\sum_{k=1}^{K}\log q_{k}$ (22)
Using the softmax function, $q_{k}=\frac{e^{z_{k}}}{\sum_{j=1}^{K}e^{z_{j}}}$, we can express $\log q_{k}$ in terms of logits:
$\log q_{k}=z_{k}-\log(\sum_{j=1}^{K}e^{z_{j}})$ (23)
Substituting this into our expression:
$\begin{split}H(\mathbf{s},\mathbf{q})=&-(1-\alpha)[z_{gt}-\log(\sum_{j=1}^{K}e^{z_{j}})]\\ &-\frac{\alpha}{K}\sum_{k=1}^{K}[z_{k}-\log(\sum_{j=1}^{K}e^{z_{j}})]\\ =&-(1-\alpha)z_{gt}+(1-\alpha)\log(\sum_{j=1}^{K}e^{z_{j}})\\ &-\frac{\alpha}{K}\sum_{k=1}^{K}z_{k}+\alpha\log(\sum_{j=1}^{K}e^{z_{j}})\\ =&-(1-\alpha)z_{gt}-\frac{\alpha}{K}\sum_{k=1}^{K}z_{k}+\log(\sum_{j=1}^{K}e^{z_{j}})\end{split}$ (24)
Rearranging:
$H(\mathbf{s},\mathbf{q})=-z_{gt}+\log(\sum_{j=1}^{K}e^{z_{j}})+\alpha[z_{gt}-\frac{1}{K}\sum_{k=1}^{K}z_{k}]$ (25)
We can identify:
  • $H(\mathbf{y},\mathbf{q})=-z_{gt}+\log(\sum_{j=1}^{K}e^{z_{j}})$ (cross-entropy for one-hot vector $\mathbf{y}$)
  • $L_{\textit{LS}}=\alpha[z_{gt}-\frac{1}{K}\sum_{k=1}^{K}z_{k}]$
Thus, we have proven:
$H(\mathbf{s},\mathbf{q})=H(\mathbf{y},\mathbf{q})+L_{\textit{LS}}$ (26)
Due to the broad usage of CutMix and Mixup in the training recipe of modern Neural Networks, we additionally take their impact into account together with Label Smoothing. Now we additionally prove the case with Cutmix and Mixup:
$L^{\prime}_{\textit{LS}}=\alpha(\left(\lambda z_{gt1}+(1-\lambda)z_{gt2}\right)-\frac{1}{K}\sum_{k=1}^{K}z_{k})$ (27)
With Cutmix and Mixup, the smoothed label becomes:
$s_{k}=(1-\alpha)(\lambda y_{k1}+(1-\lambda)y_{k2})+\frac{\alpha}{K}$ (28)
where $y_{k1}$ and $y_{k2}$ are one-hot vectors for the two ground truth classes from mixing, and $\lambda$ is the mixing ratio.
Starting with the cross-entropy:
$H(\mathbf{s},\mathbf{q})$ $=-\sum_{k=1}^{K}s_{k}\log q_{k}$ (29)
$=-\sum_{k=1}^{K}[(1-\alpha)(\lambda y_{k1}+(1-\lambda)y_{k2})+\frac{\alpha}{K}]\log q_{k}$ (30)
$=-(1-\alpha)\sum_{k=1}^{K}(\lambda y_{k1}+(1-\lambda)y_{k2})\log q_{k}-\frac{\alpha}{K}\sum_{k=1}^{K}\log q_{k}$ (31)
Since $y_{k1}$ and $y_{k2}$ are one-hot vectors:
$H(\mathbf{s},\mathbf{q})=-(1-\alpha)(\lambda\log q_{gt1}+(1-\lambda)\log q_{gt2})-\frac{\alpha}{K}\sum_{k=1}^{K}\log q_{k}$ (32)
where $gt1$ and $gt2$ are the indices of the two ground truth classes.
Using $q_{k}=\frac{e^{z_{k}}}{\sum_{j=1}^{K}e^{z_{j}}}$, we express in terms of logits:
$H(\mathbf{s},\mathbf{q})$ $=-(1-\alpha)[\lambda(z_{gt1}-\log(\sum_{j=1}^{K}e^{z_{j}}))+(1-\lambda)(z_{gt2}-\log(\sum_{j=1}^{K}e^{z_{j}}))]$ (33)
$\quad-\frac{\alpha}{K}\sum_{k=1}^{K}[z_{k}-\log(\sum_{j=1}^{K}e^{z_{j}})]$ (34)
Simplifying:
$H(\mathbf{s},\mathbf{q})$ $=-(1-\alpha)[\lambda z_{gt1}+(1-\lambda)z_{gt2}]+(1-\alpha)\log(\sum_{j=1}^{K}e^{z_{j}})$ (35)
$\quad-\frac{\alpha}{K}\sum_{k=1}^{K}z_{k}+\alpha\log(\sum_{j=1}^{K}e^{z_{j}})$ (36)
$=-(1-\alpha)[\lambda z_{gt1}+(1-\lambda)z_{gt2}]-\frac{\alpha}{K}\sum_{k=1}^{K}z_{k}+\log(\sum_{j=1}^{K}e^{z_{j}})$ (37)
Rearranging:
$H(\mathbf{s},\mathbf{q})$ $=-[\lambda z_{gt1}+(1-\lambda)z_{gt2}]+\log(\sum_{j=1}^{K}e^{z_{j}})$ (38)
$\quad+\alpha[\lambda z_{gt1}+(1-\lambda)z_{gt2}-\frac{1}{K}\sum_{k=1}^{K}z_{k}]$ (39)
We can identify:
  • $H(\mathbf{y}^{\prime},\mathbf{q})=-[\lambda z_{gt1}+(1-\lambda)z_{gt2}]+\log(\sum_{j=1}^{K}e^{z_{j}})$ (cross-entropy for mixed label $\mathbf{y}^{\prime}$)
  • $L^{\prime}_{\textit{LS}}=\alpha[\lambda z_{gt1}+(1-\lambda)z_{gt2}-\frac{1}{K}\sum_{k=1}^{K}z_{k}]$
Thus, we have proven:
$H(\mathbf{s},\mathbf{q})=H(\mathbf{y}^{\prime},\mathbf{q})+L^{\prime}_{\textit{LS}}$ (40)
This completes the proof for both cases of Theorem 3.3.
Appendix C Gradient Analysis
C.1 New Objective Function
The Cross Entropy with Max Suppression is defined as:
$L_{\text{MaxSup},t}(x,y)=H\left(y_{k}+\frac{\alpha}{K}-\alpha\cdot\mathbf{1}_{k=\operatorname{argmax}({\bm{q}})},{\bm{q}}_{t}^{S}(x)\right)$
where $H(\cdot,\cdot)$ denotes the cross-entropy function.
C.2 Gradient Analysis
The gradient of the loss with respect to the logit $z_{i}$ for each class $i$ is derived as:
$\partial_{i}^{\text{MaxSup},t}=y_{t,i}-y_{i}-\frac{\alpha}{K}+\alpha\cdot\mathbf{1}_{i=\operatorname{argmax}({\bm{q}})}$
We analyze this gradient under two scenarios:
Scenario 1: Model makes correct prediction
In this case, Max Suppression is equivalent to Label Smoothing. When the model correctly predicts the target class ($\operatorname{argmax}({\bm{q}})=\operatorname{GT}$), the gradients are:
  • For the target class (GT): $\partial_{\text{GT}}^{\text{MaxSup},t}=q_{t,\text{GT}}-\left(1-\alpha\left(1-\frac{1}{K}\right)\right)$
  • For non-target classes: $\partial_{i}^{\text{MaxSup},t}=q_{t,i}-\frac{\alpha}{K}$
Scenario 2: Model makes wrong prediction
When the model incorrectly predicts the most confident class ($\operatorname{argmax}({\bm{q}})\neq\operatorname{GT}$), the gradients are:
  • For the target class (GT): $\partial_{\text{GT}}^{\text{MaxSup},t}=q_{t,\text{GT}}-\left(1+\frac{\alpha}{K}\right)$
  • For non-target classes (not most confident): $\partial_{i}^{\text{MaxSup},t}=q_{t,i}-\frac{\alpha}{K}$
  • For the most confident non-target class: $\partial_{i}^{\text{MaxSup},t}=q_{t,i}+\alpha\left(1-\frac{1}{K}\right)$
The Max Suppression regularization technique implements a sophisticated gradient redistribution strategy, particularly effective when the model misclassifies samples. When the model’s prediction ($\operatorname{argmax}({\bm{q}})$) differs from the ground truth (GT), the gradient for the incorrectly predicted class is increased by $\alpha(1-\frac{1}{K})$, resulting in $\partial_{\operatorname{argmax}({\bm{q}})}^{\text{MaxSup},t}=q_{t,\operatorname{argmax}({\bm{q}})}+\alpha(1-\frac{1}{K})$. Simultaneously, the gradient for the true class is decreased by $\frac{\alpha}{K}$, giving $\partial_{\text{GT}}^{\text{MaxSup},t}=q_{t,\text{GT}}-(1+\frac{\alpha}{K})$, while for all other classes, the gradient is slightly reduced by $\frac{\alpha}{K}$: $\partial_{i}^{\text{MaxSup},t}=q_{t,i}-\frac{\alpha}{K}$. This redistribution adds a substantial positive gradient to the misclassified class while slightly reducing the gradients for other classes. The magnitude of this adjustment, controlled by the hyperparameter $\alpha$, effectively penalizes overconfident errors and encourages the model to focus on challenging examples. By amplifying the learning signal for misclassifications, Max Suppression regularization promotes more robust learning from difficult or ambiguous samples.
Algorithm 1 Gradient Descent with Max Suppression (MaxSup)
[…]
$L_{\mathrm{CE}}\leftarrow\frac{1}{B}\sum_{j=1}^{B}H\bigl{(}\mathbf{y}^{(j)},\mathbf{q}^{(j)}\bigr{)}$
   // MaxSup component: penalize the top-1 logitFor each sample $j$:
$z_{\textit{max}}^{(j)}\;=\;\max_{k\in\{1,\dots,K\}}z_{k}^{(j)},\quad\bar{z}^{(j)}\;=\;\frac{1}{K}\sum_{k=1}^{K}z_{k}^{(j)}$
$L_{\mathrm{MaxSup}}\leftarrow\frac{1}{B}\sum_{j=1}^{B}\bigl{[}z_{\textit{max}}^{(j)}-\bar{z}^{(j)}\bigr{]}$
Total loss:
$L\;\leftarrow\;L_{\mathrm{CE}}\;+\;\alpha\,L_{\mathrm{MaxSup}}$
Update parameters:
$\theta\;\leftarrow\;\theta-\eta\,\nabla_{\theta}\,L$
[…]
Appendix D Pseudo Code
Algorithm 1 presents pseudo code illustrating gradient descent with Max Suppression (MaxSup). The main difference from standard Label Smoothing lies in penalizing the highest logit rather than the ground-truth logit.
Appendix E Robustness Under Different Training Recipes
We assess MaxSup’s robustness by testing it under a modified training recipe that reduces total training time and alters the learning rate schedule. This setup models scenarios where extensive training is impractical due to limited resources.
Concretely, we adopt the TorchVision V1 Weight strategy, reducing the total number of epochs to 90 and replacing the cosine annealing schedule with a step learning-rate scheduler (step size = 30). We also set the initial learning rate to 0.1 and use a batch size of 512. This streamlined recipe aims to reach reasonable accuracy within a shorter duration.
As reported in Table 7, MaxSup continues to deliver strong performance across multiple convolutional architectures, generally surpassing Label Smoothing and its variants. Although all methods see a performance decline in this constrained regime, MaxSup remains among the top performers, reinforcing its effectiveness across diverse training conditions.
Table 7: Performance comparison on ImageNet for various convolutional neural network architectures. Results are presented as “mean ± std” (percentage). Bold and underlined entries indicate best and second-best, respectively. (: implementation details adapted from the official repositories.)
Method ResNet-18 ResNet-50 ResNet-101 MobileNetV2
Baseline 69.11±0.12 76.44±0.10 76.00±0.18 71.42±0.12
Label Smoothing 69.38±0.19 76.65±0.11 77.01±0.15 71.40±0.09
Zipf-LS 69.43±0.13 76.89±0.17 76.91±0.14 71.24±0.16
OLS 69.45±0.15 76.81±0.21 77.12±0.17 71.29±0.11
MaxSup 69.59±0.13 77.08±0.07 77.33±0.12 71.59±0.17
Logit Penalty 66.97±0.11 74.21±0.16 75.17±0.12 70.39±0.14
Appendix F Increasing Smoothing Weight Schedule
Building on the intuition that a model’s confidence naturally grows as training progresses, we propose a linearly increasing schedule for the smoothing parameter $\alpha$. Concretely, $\alpha$ is gradually raised from an initial value (e.g., 0.1) to a higher value (e.g., 0.2) by the end of training. This schedule aims to counteract the model’s increasing overconfidence, ensuring that regularization remains appropriately scaled throughout.
Experimental Evidence
As shown in Table 8, both Label Smoothing and MaxSup benefit from this $\alpha$ scheduler. For Label Smoothing, accuracy improves from 75.91% to 76.16%, while MaxSup sees a more pronounced gain, from 76.12% to 76.58%. This greater improvement for MaxSup ($+0.46\%$) compared to Label Smoothing ($+0.25\%$) corroborates our claim that MaxSup successfully addresses the inconsistent regularization and error-enhancement issues of Label Smoothing during misclassifications.
Table 8: Effect of an $\alpha$ scheduler on model performance. Here, $t$ and $T$ denote the current and total epochs, respectively. The baseline model does not involve any label smoothing parameter ($\alpha$).
Configuration Formulation $\alpha=0.1$ $\alpha=0.1+0.1\,\tfrac{t}{T}$ Remarks
Baseline 74.21 74.21 $\alpha$ not used
LS $\alpha\Bigl{(}z_{gt}-\tfrac{1}{K}\sum_{k}z_{k}\Bigr{)}$ 75.91 76.16
MaxSup $\alpha\Bigl{(}z_{\textit{max}}-\tfrac{1}{K}\sum_{k}z_{k}\Bigr{)}$ 76.12 76.58
Appendix G Visualization of the Learned Feature Space
To illustrate the differences between Max Suppression Regularization and Label Smoothing, we follow the projection technique of Müller et al. (2019). Specifically, we select three semantically related classes and construct an orthonormal basis for the plane intersecting their class templates in feature space. We then project each sample’s penultimate-layer activation vector onto this plane. To ensure the visual clarity of the resulting plots, we randomly sample 80 images from the training or validation set for each of the three classes.
Selection Criteria
We choose these classes according to two main considerations:
  1. 1.
    Semantic Similarity. We pick three classes that are visually and semantically close.
  2. 2.
    Confusion. We identify a class that the Label Smoothing (LS)–trained model frequently misclassifies and select two additional classes involved in those misclassifications (Figure 4(c), Figure 5(c)). Conversely, we also examine a scenario where a class under Max Suppression is confused with others, highlighting key differences (Figure 4(d), Figure 5(d)).
(a) Semantically Similar Classes
(b) Semantically Similar Classes
(c) Confusing Classes (LS)
(d) Confusing Classes (MaxSup)
Figure 4: Visualization of penultimate-layer activations from DeiT-Small (trained with CutMix and Mixup) on the ImageNet validation set. The top row shows embeddings for a MaxSup-trained model, and the bottom row shows embeddings for a Label Smoothing (LS)–trained model. In each subfigure, classes are either semantically similar or confusingly labeled. Compared to LS, MaxSup yields more pronounced inter-class separability and richer intra-class diversity, suggesting stronger representation and classification performance.
(a) Semantically Similar Classes
(b) Semantically Similar Classes
(c) Confusing Classes (LS)
(d) Confusing Classes (MaxSup)
Figure 5: Visualization of the penultimate-layer activations for DeiT-Small (trained with CutMix and Mixup) on selected ImageNet classes. The top row shows results for a MaxSup-trained model; the bottom row shows Label Smoothing (LS). In (a,b), the model must distinguish semantically similar classes (e.g., Saluki vs. Grey Fox; Tow Truck vs. Pickup), while (c,d) involve confusing categories (e.g., Jean vs. Shoe Shop, Stinkhorn vs. related objects). Compared to LS, MaxSup yields both improved inter-class separability and richer intra-class variation, indicating more robust representation learning.
Observations
As shown in Figures 5 and 4, models trained with Max Suppression exhibit:
  • Enhanced inter-class separability. Distinct classes occupy more clearly separated regions, aligning with improved classification performance.
  • Greater intra-class variation. Instances within a single class are not overly compressed, indicating a richer representation of subtle differences.
For instance, images of Schipperke dogs can differ markedly in viewpoint, lighting, background, or partial occlusions. Max Suppression preserves such intra-class nuances in the feature space, enabling the semantic distances to visually related classes (e.g., Saluki, Grey Fox, or Belgian Sheepdog) to dynamically adjust for each image. Consequently, Max Suppression provides a more flexible, fine-grained representation that facilitates better class discrimination.
Table 9: Feature representation metrics for a ResNet-50 model trained on ImageNet-1K, reported on both Training and Validation sets. We measure intra-class variation ($\bar{d}_{\text{within}}$) and overall average distance ($\bar{d}_{\text{total}}$). Inter-class separability ($R^{2}$) is calculated as $R^{2}=1-\frac{\bar{d}_{\text{within}}}{\bar{d}_{\text{total}}}$. Higher values ($\uparrow$) of $\bar{d}_{\text{within}}$ and $R^{2}$ are preferred.
Method $\bar{d}_{\text{within}}\uparrow$ $\bar{d}_{\text{total}}$ $R^{2}\uparrow$
Train Val Train Val Train Val
Baseline 0.3114 0.3313 0.5212 0.5949 0.4025 0.4451
LS 0.2632 0.2543 0.4862 0.4718 0.4690 0.4611
OLS 0.2707 0.2820 0.6672 0.6570 0.5943 0.5708
Zipf’s 0.2611 0.2932 0.5813 0.5628 0.5522 0.4790
MaxSup 0.2926 (+0.03) 0.2998 (+0.05) 0.6081 (+0.12) 0.5962 (+0.12) 0.5188 (+0.05) 0.4972 (+0.04)
Logit Penalty 0.2840 0.3144 0.7996 0.7909 0.6448 0.6024

### v2

@@
 Abstract
-Label Smoothing (LS) is widely adopted to curb overconfidence in neural network predictions and enhance generalization.
-However, previous research shows that LS can force feature representations into excessively tight clusters, eroding intra-class distinctions.
-More recent findings suggest that LS also induces overconfidence in misclassifications, yet the precise mechanism remained unclear.
-In this work, we decompose the loss term introduced by LS, revealing two key components: (i) a regularization term that functions only when the prediction is correct, and (ii) an error-enhancement term that emerges under misclassifications.
-This latter term compels the model to reinforce incorrect predictions with exaggerated certainty, further collapsing the feature space.
-To address these issues, we propose Max Suppression (MaxSup), which uniformly applies the intended regularization to both correct and incorrect predictions by penalizing the top-1 logit instead of the ground-truth logit.
-Through feature analyses, we show that MaxSup restores intra-class variation and sharpens inter-class boundaries.
-Extensive experiments on image classification and downstream tasks confirm that MaxSup is a more robust alternative to LS.
-[^1] Code is available at: https://github.com/ZhouYuxuanYX/Maximum-Suppression-Regularization.
-Machine Learning, ICML
+Label Smoothing (LS) is widely adopted to reduce overconfidence in neural network predictions and improve generalization.
+Despite these benefits, recent studies reveal two critical issues with LS.
+First, LS induces overconfidence in misclassified samples.
+Second, it compacts feature representations into overly tight clusters, diluting intra-class diversity, although the precise cause of this phenomenon remained elusive.
+In this paper, we analytically decompose the LS-induced loss, exposing two key terms: (i) a regularization term that dampens overconfidence only when the prediction is correct, and (ii) an error-amplification term that arises under misclassifications.
+This latter term compels the network to reinforce incorrect predictions with undue certainty, exacerbating representation collapse.
+To address these shortcomings, we propose Max Suppression (MaxSup), which applies uniform regularization to both correct and incorrect predictions by penalizing the top-1 logit rather than the ground-truth logit.
+Through extensive feature-space analyses, we show that MaxSup restores intra-class variation and sharpens inter-class boundaries.
+Experiments on large-scale image classification and multiple downstream tasks confirm that MaxSup is a more robust alternative to LS, consistently reducing overconfidence while preserving richer feature representations.
+[^1] Code is available at: https://anonymous.4open.science/r/Maximum-Suppression-Regularization-DB0C.
+1 Introduction
+Multi-class classification [17, 24] typically relies on one-hot labels, which implicitly treat different classes as mutually orthogonal.
+In practice, however, classes often share low-level features [29, 40] or exhibit high-level semantic similarities [3, 22, 38], rendering the one-hot assumption overly restrictive.
+Such a mismatch can yield over-confident classifiers and ultimately degrade generalization [10].
+To moderate overconfidence, Szegedy et al.
+[32] introduced Label Smoothing (LS), which combines a uniform distribution with the hard ground-truth label, thereby reducing the model’s certainty in the primary class.
+LS has since become prevalent in image recognition [11, 20, 33, 43] and neural machine translation [1, 6], often boosting accuracy and calibration [21].
+Yet subsequent work indicates that LS can overly compress features into tight clusters [13, 26, 37], hindering intra-class variability and transferability [5].
+In parallel, Zhu et al.
+[44] found that LS paradoxically fosters overconfidence in misclassified samples, though the precise mechanism behind this remains uncertain.
+In this paper, we reveal that LS’s training objective inherently contains an error amplification term.
+This term pushes the network to reinforce incorrect predictions with exaggerated certainty, yielding highly confident misclassifications and further compressing feature clusters (Section 3.1, Table 1).
+Building on Zhu et al.
+[44], we characterize “overconfidence” in terms of the model’s top-1 prediction, rather than through conventional calibration metrics.
+Through our analysis, we further show that punishing the ground-truth logit during misclassification reduces intra-class variation (Table 2), a phenomenon corroborated by Grad-CAM visualizations (Figure 2).
+To overcome these shortcomings, we introduce Max Suppression (MaxSup), a method that retains the beneficial regularization effect of LS while eliminating its error amplification.
+Rather than penalizing the ground-truth logit, MaxSup focuses on the model’s top-1 logit, ensuring a consistent regularization signal regardless of whether the current prediction is correct or misclassified.
+By preserving the ground-truth logit in misclassifications, MaxSup sustains richer intra-class variability and sharpens inter-class boundaries.
+As visualized in Figure 1, this approach mitigates the feature collapse and attention drift often induced by LS, ultimately leading to more robust representations.
+Through comprehensive experiments in both image classification (Section 4.2) and semantic segmentation (Section 4.3), we show that MaxSup not only alleviates severe intra-class collapse but also consistently boosts top-1 accuracy (+0.4%–1.0% improvement over LS on ImageNet) and robustly enhances downstream transfer performance (Section 4.1).
+Our contributions are summarized as follows:
+  • We perform a logit-level analysis of Label Smoothing, revealing how the error amplification term inflates misclassification confidence and compresses features.
+  • We propose Max Suppression (MaxSup), removing detrimental error amplification while preserving LS’s beneficial regularization.
+As shown in extensive ablations, MaxSup alleviates intra-class collapse and yields consistent accuracy gains.
+  • We demonstrate superior performance across tasks and architectures, including ResNet, MobileNetV2, and DeiT-S, where MaxSup significantly boosts accuracy on ImageNet (up to +1.0% in top-1) and consistently delivers stronger representations for downstream tasks such as semantic segmentation and robust transfer learning.
 Figure 1: Comparison of Label Smoothing (LS) and MaxSup.
@@
 Right: Grad-CAM visualizations show that MaxSup more effectively highlights class-discriminative regions than LS.
-1 Introduction
-Multi-class classification (Russakovsky et al., 2015; LeCun, 1998) conventionally relies on one-hot labels, implicitly treating each class as if it were completely orthogonal to every other.
-In reality, however, classes often share low-level attributes (Zeiler & Fergus, 2014; Silla & Freitas, 2011) or exhibit high-level semantic similarities (Chen et al., 2021; Yi et al., 2022; Novack et al., 2023), which makes the strict orthogonality assumption overly simplistic.
-This mismatch can lead to over-confident classifiers, ultimately reducing generalization (Guo et al., 2020).
-To address overconfidence, Szegedy et al.
-(2016) introduced Label Smoothing (LS), blending a uniform distribution with the hard label to reduce the model’s certainty in the target class.
-LS has become a mainstay in both image recognition (He et al., 2016; Touvron et al., 2021; Liu et al., 2021; Zhou et al., 2022b) and neural machine translation (Gao et al., 2020; Alves et al., 2023), often improving accuracy and calibration (Müller et al., 2019).
-However, studies have also revealed that LS can produce overly tight clusters in the feature space (Kornblith et al., 2021; Sariyildiz et al., 2022; Xu & Liu, 2023), thereby lowering intra-class diversity and harming transferability (Feng et al., 2021).
-Meanwhile, Zhu et al.
-(2022) reported that LS inadvertently increases confidence in incorrect predictions, though the exact cause remained unclear.
-In this paper, we show that LS’s training objective inherently contains an error-enhancement term that amplifies misclassified predictions, thus causing overconfident errors and tighter feature clusters (Section 3.1, Table 1).
-Extending Zhu et al.
-(2022), we define “overconfidence” in terms of the network’s top-1 prediction rather than calibration-based criteria.
-Our analysis further demonstrates that penalizing the ground-truth logit in misclassifications compresses the feature space, reducing intra-class variation (Table 2), as corroborated by Grad-CAM visualizations (Figure 3).
-To overcome these limitations, we propose Max Suppression (MaxSup), which retains LS’s desirable regularization effect while eliminating its error-enhancement component.
-Rather than penalizing the ground-truth logit, MaxSup penalizes the largest logit, thus providing consistent regularization regardless of prediction correctness.
-By preventing ground-truth suppression during misclassification, MaxSup preserves richer intra-class variation and improves inter-class separability.
-As illustrated in Figure 1, this alleviates the compression and attentional shortcomings introduced by LS, leading to more robust feature representations.
-Extensive experiments on both image classification and semantic segmentation confirm that MaxSup not only alleviates intra-class collapse but also boosts final accuracy, enhances generalization, and strengthens transfer performance.
-Our contributions are summarized as follows:
-  • We present a logit-level analysis of Label Smoothing that unearths an ‘error-enhancement” term, revealing how LS inadvertently reinforces overconfidence in misclassified samples.
-  • We propose Max Suppression (MaxSup), which preserves LS’s desired regularization while removing its detrimental error-enhancement component, thereby reducing intra-class compression and boosting both classification and downstream task performance.
 2 Related Work
-We survey regularization techniques before focusing on Label Smoothing (LS) and highlighting how MaxSup differs.
+We first outline mainstream regularization techniques in deep learning, then survey recent advances in Label Smoothing (LS), and finally clarify how our MaxSup diverges from prior variants.
 2.1 Regularization
-Regularization enhances the generalization of deep neural networks by limiting model complexity through various strategies.
-Classical approaches such as $\ell_{2}$ (Krogh & Hertz, 1991) and $\ell_{1}$ (Zou & Hastie, 2005) regularization constrain large or sparse weights, respectively, while Dropout (Srivastava et al., 2014) randomly deactivates neurons to prevent feature co-adaptation.
-Among loss-based methods, Label Smoothing (LS) (Szegedy et al., 2016) redistributes probability mass away from the ground-truth class, improving both accuracy and calibration (Müller et al., 2019).
-Variants like Online Label Smoothing (OLS) (Zhang et al., 2021) and Zipf Label Smoothing (Zipf-LS) (Liang et al., 2022) adapt LS by considering the model’s evolving predictions, yet they still fail to address the fundamental issue that arises when the ground-truth logit is not maximal (Section 3.1, Table 1).
-Other loss-based regularizers, such as Confidence Penalty (Pereyra et al., 2017) and Logit Penalty (Dauphin & Cubuk, 2021), target different aspects of the output distribution.
-Confidence Penalty discourages overconfident predictions, whereas Logit Penalty minimizes the global $\ell_{2}$-norm of logits to improve feature separability (Kornblith et al., 2021).
-However, Logit Penalty can reduce intra-class variation, impairing transfer learning (Section 4).
-Our MaxSup approach
-MaxSup diverges from these methods by selectively penalizing only the top-1 logit $(z_{\textit{max}})$ rather than the ground-truth logit $(z_{gt})$.
-Unlike LS-based techniques, which can exacerbate errors by excessively shrinking $z_{gt}$ for misclassified samples, MaxSup uniformly applies regularization to all predictions, regardless of correctness.
-Consequently, it effectively sidesteps the error-enhancement issue, preserves richer intra-class diversity (Table 2), and sustains robust transfer performance across various datasets and architectures (Table 3).
+Regularization techniques aim to improve the generalization of deep neural networks by constraining model complexity.
+Classical methods like $\ell_{2}$ [16] and $\ell_{1}$ [45] impose direct penalties on large or sparse weights, while Dropout [30] randomly deactivates neurons to discourage over-adaptation.
+In the realm of loss-based strategies, Label Smoothing (LS) [32] redistributes a fraction of the label probability mass away from the ground-truth class, thereby improving accuracy and calibration [21].
+Variants such as Online Label Smoothing (OLS) [41] and Zipf Label Smoothing (Zipf-LS) [19] refine LS by dynamically adjusting the smoothed labels based on a model’s evolving predictions.
+However, they do not fully address the fundamental issue that emerges when the ground-truth logit is not the highest one (see Section 3.1, Table 1).
+Other loss-based regularizers focus on alternative aspects of the predictive distribution.
+Confidence Penalty [23] penalizes the model’s confidence directly, while Logit Penalty [4] minimizes the global $\ell_{2}$-norm of logits, a technique reported to enhance class separation [13].
+Despite these benefits, Logit Penalty can inadvertently shrink intra-class variation, thereby hampering transfer learning (see Section 4.1).
+Unlike the aforementioned methods, MaxSup enforces regularization by penalizing only the top-1 logit $z_{\textit{max}}$ rather than the ground-truth logit $z_{gt}$.
+In LS-based approaches, suppressing $z_{gt}$ for misclassified samples can worsen errors, whereas MaxSup applies a uniform penalty regardless of whether the model’s prediction is correct.
+Consequently, MaxSup avoids the error amplification effect, retains richer intra-class diversity (see Table 2), and achieves robust transfer performance across diverse datasets and model families (see Table 3).
 2.2 Studies on Label Smoothing
-Label Smoothing (LS) has also been extensively examined in the context of knowledge distillation.
-Yuan et al.
-(2020) showed that LS can act as a proxy for distillation, while Shen et al.
-(2021) explored its role within teacher–student frameworks.
-Chandrasegaran et al.
-(2022) further demonstrated that a low-temperature, LS-trained teacher can improve distillation performance.
-Meanwhile, Kornblith et al.
-(2021) found that LS tightens class clusters in feature space, reducing transfer performance.
-From a Neural Collapse (NC) perspective (Zhou et al., 2022a; Guo et al., 2024), LS drives the model toward rigid feature clusters, a phenomenon measured by Xu & Liu (2023) via a variability metric.
-Comparison with existing LS techniques
-Our primary objective is to mitigate the error-enhancement effect.
-Instead of refining a smoothed label, as in OLS or Zipf-LS, MaxSup directly penalizes the highest logit $z_{\textit{max}}$.
-This simple yet effective modification ensures uniform regularization even when $z_{gt}$ is not the top logit, thereby maintaining greater intra-class diversity and avoiding the performance degradation common to LS-based approaches (Section 3.2).
-Additionally, MaxSup integrates seamlessly with standard training pipelines, requiring no extra computational overhead beyond simply replacing LS.
+Label Smoothing has also been studied extensively under knowledge distillation.
+For instance, Yuan et al.
+[39] observed that LS can approximate the effect of a teacher–student framework, while Shen et al.
+[28] investigated its role in such pipelines more systematically.
+Additionally, Chandrasegaran et al.
+[2] demonstrated that a low-temperature, LS-trained teacher can notably improve distillation outcomes.
+Concurrently, Kornblith et al.
+[13] showed that LS tightens intra-class clusters in the feature space, diminishing transfer performance.
+From a Neural Collapse perspective [42, 9], LS nudges the model toward rigid feature clusters, as evidenced by the reduced feature variability measured in Xu and Liu [37].
+Our goal is to overcome LS’s inherent error amplification effect.
+Rather than adjusting how the smoothed label distribution is constructed (as in OLS or Zipf-LS), MaxSup directly penalizes the highest logit $z_{\textit{max}}$.
+This design ensures consistent regularization even if $z_{gt}$ is not the top logit, thereby avoiding the degradation in performance typical of misclassified samples under LS (see Section 3.2).
+Moreover, MaxSup integrates seamlessly into standard training pipelines, introducing negligible computational overhead beyond substituting the LS term.
 3 Max Suppression Regularization (MaxSup)
 We first partition the training objective into two components: the standard Cross-Entropy (CE) loss and a regularization term introduced by Label Smoothing (LS).
-By expressing LS in terms of logits (Theorem 3.3), we isolate two key factors: a regularization term that controls overconfidence and an error-enhancement term that enlarges the gap between the ground-truth logit $z_{gt}$ and any higher logits (Corollary 3.4, Equation 5), ultimately degrading performance.
-To address these shortcomings, we propose Max Suppression Regularization (MaxSup), which applies the penalty to the largest logit $z_{\textit{max}}$ rather than to $z_{gt}$ (Equation 8, Section 3.2).
+By expressing LS in terms of logits (Theorem 3.3), we isolate two key factors: a regularization term that controls overconfidence and an error amplification term that enlarges the gap between the ground-truth logit $z_{gt}$ and any higher logits (Corollary 3.4, Equation 5), ultimately degrading performance.
+To address these issues, we propose Max Suppression Regularization (MaxSup), which applies the penalty to the largest logit $z_{\textit{max}}$ rather than $z_{gt}$ (Equation 8, Section 3.2).
 This shift delivers consistent regularization for both correct and incorrect predictions, preserves intra-class variation, and bolsters inter-class separability.
 Consequently, MaxSup mitigates the representation collapse found in LS, attains superior ImageNet-1K accuracy (Table 1), and improves transferability (Table 2, Table 3).
-The following sections elaborate on MaxSup’s formulation and its integration into the overall training pipeline.
+The following sections elaborate on MaxSup’s formulation and integration into the training pipeline.
 3.1 Revisiting Label Smoothing
@@
 Definition 3.1.
-For a classification task with $K$ classes, LS converts a one-hot label $\mathbf{y}\in\mathbb{R}^{K}$ into a soft label $\mathbf{s}\in\mathbb{R}^{K}$:
+For a standard classification task with $K$ classes, Label Smoothing (LS) converts a one-hot label $\mathbf{y}\in\mathbb{R}^{K}$ into a softened target label $\mathbf{s}\in\mathbb{R}^{K}$:
 $s_{k}=(1-\alpha)y_{k}+\frac{\alpha}{K},$ (1)
@@
 $L_{\textit{LS}}\;=\;\alpha\,\biggl{(}H\Bigl{(}\tfrac{\mathbf{1}}{K},\mathbf{q}\Bigr{)}\;-\;H(\mathbf{y},\mathbf{q})\biggr{)}.$ (3)
-Here, $\mathbf{q}$ is the predicted probability vector, $H(\cdot)$ denotes the Cross-Entropy, and $\frac{\mathbf{1}}{K}$ is the uniform distribution introduced by LS.
+Where, $\mathbf{q}$ is the predicted probability vector, $H(\cdot)$ denotes the Cross-Entropy, and $\frac{\mathbf{1}}{K}$ is the uniform distribution introduced by LS.
 This shows that LS adds a regularization term, $L_{\textit{LS}}$, which smooths the output distribution and helps to reduce overfitting.
@@
 Decomposition of Label Smoothing Loss.
-$L_{\textit{LS}}\;=\;\underbrace{\frac{\alpha}{K}\sum_{z_{m}<z_{gt}}\Bigl{(}z_{gt}-z_{m}\Bigr{)}}_{\text{Regularization}}\;+\;\underbrace{\frac{\alpha}{K}\sum_{z_{n}>z_{gt}}\Bigl{(}z_{gt}-z_{n}\Bigr{)}}_{\text{Error-Enhancement}},$ (5)
-where $M$ and $N$ are the numbers of logits below and above $z_{gt}$, respectively ($M+N=K-1$). Note that the error-enhancement term vanishes when $z_{gt}=z_{\textit{max}}$.
+$L_{\textit{LS}}\;=\;\underbrace{\frac{\alpha}{K}\sum_{z_{m}<z_{gt}}\Bigl{(}z_{gt}-z_{m}\Bigr{)}}_{\text{Regularization}}\;+\;\underbrace{\frac{\alpha}{K}\sum_{z_{n}>z_{gt}}\Bigl{(}z_{gt}-z_{n}\Bigr{)}}_{\text{error amplification}},$ (5)
+where $M$ and $N$ are the numbers of logits below and above $z_{gt}$, respectively ($M+N=K-1$). Note that the error amplification term vanishes when $z_{gt}=z_{\textit{max}}$.
   1. 1.
@@
   2. 2.
-    Error-Enhancement: Penalizes the gap between $z_{gt}$ and larger logits, inadvertently increasing overconfidence in incorrect predictions.
-Although LS aims to combat overfitting by reducing prediction confidence, its error-enhancement component can be detrimental for misclassified samples, as it widens the gap between the ground-truth logit $z_{gt}$ and the incorrect top logit.
+    Error amplification: Penalizes the gap between $z_{gt}$ and larger logits, inadvertently increasing overconfidence in incorrect predictions.
+Although LS aims to combat overfitting by reducing prediction confidence, its error amplification component can be detrimental for misclassified samples, as it widens the gap between the ground-truth logit $z_{gt}$ and the incorrect top logit.
 Concretely:
   1. 1.
-    Correct Predictions $(z_{gt}=z_{\textit{max}})$: The error-enhancement term is zero, and the regularization term effectively reduces overconfidence by shrinking the gap between $z_{gt}$ and any smaller logits.
+    Correct Predictions $(z_{gt}=z_{\textit{max}})$: The error amplification term is zero, and the regularization term effectively reduces overconfidence by shrinking the gap between $z_{gt}$ and any smaller logits.
   2. 2.
     Incorrect Predictions $(z_{gt}\neq z_{\textit{max}})$: LS introduces two potential issues:
-    • Error-Enhancement: Increases the gap between $z_{gt}$ and larger logits, reinforcing overconfidence in incorrect predictions.
+    • Error amplification: Increases the gap between $z_{gt}$ and larger logits, reinforcing overconfidence in incorrect predictions.
     • Inconsistent Regularization: The regularization term lowers $z_{gt}$ yet does not penalize $z_{\textit{max}}$, which further impairs learning.
-These issues with LS on misclassified samples have also been observed in prior work (Xia et al., 2024).
-By precisely identifying these two components (regularization vs.
-error-enhancement), we can design a more targeted solution.
-Ablation Study on LS Components
-To gauge the influence of each LS component, we conduct an ablation study on ImageNet-1K using a DeiT-Small model (Touvron et al., 2021) without Mixup or CutMix.
-As shown in Table 1, LS’s gains arise solely from the regularization term, whereas the error-enhancement term degrades performance.
-In contrast, our proposed Max Suppression Regularization (MaxSup) omits the error-enhancement effect and achieves higher accuracy by retaining the beneficial regularization.
+These issues with LS on misclassified samples have also been systematically observed in prior work [35].
+By precisely disentangling these two components (regularization vs.
+error amplification), we can design a more targeted and effective solution.
 Table 1: Ablation on LS components using DeiT-Small on ImageNet-1K (without CutMix or Mixup).
-“Regularization” denotes penalizing logits smaller than $z_{gt}$; “Error-Enhancement” penalizes logits larger than $z_{gt}$.
-MaxSup removes error-enhancement while retaining regularization.
+“Regularization” denotes penalizing logits smaller than $z_{gt}$; “error amplification” penalizes logits larger than $z_{gt}$.
+MaxSup removes error amplification while retaining regularization.
 Method Formulation Accuracy
@@
 + Regularization $\tfrac{\alpha}{M}\sum_{z_{m}<z_{gt}}(z_{gt}-z_{m})$ 75.98
-+ Error-Enhancement $\tfrac{\alpha}{N}\sum_{z_{n}>z_{gt}}(z_{gt}-z_{n})$ 73.63
-+ Error-Enhancement $\alpha\,(z_{gt}-z_{\textit{max}})$ 73.69
++ error amplification $\tfrac{\alpha}{N}\sum_{z_{n}>z_{gt}}(z_{gt}-z_{n})$ 73.63
++ error amplification $\alpha\,(z_{gt}-z_{\textit{max}})$ 73.69
 + MaxSup $\alpha\Bigl{(}z_{\textit{max}}-\tfrac{1}{K}\!\sum_{k=1}^{K}z_{k}\Bigr{)}$ 76.12
-From Table 1, it is evident that LS’s overall accuracy boost is exclusively attributed to the regularization component, whereas error-enhancement consistently degrades performance (73.63% or 73.69%).
-Removing the error-enhancement term while keeping only the regularization improves accuracy slightly (75.98% vs.
+Ablation Study on LS Components.
+To isolate the effects of each component in LS, we carefully perform a detailed and systematic ablation study on ImageNet-1K using a DeiT-Small model [33] without Mixup or CutMix.
+As indicated in Table 1, the performance gains from LS stem solely from the regularization term, whereas the error amplification term degrades accuracy.
+In contrast, our MaxSup omits the error amplification component and leverages only the beneficial regularization, thereby boosting accuracy beyond that of standard LS.
+Specifically, Table 1 shows that LS’s overall improvement can be attributed exclusively to its regularization contribution; the error amplification term consistently reduces accuracy (e.g., to 73.63% or 73.69%).
+Disabling only the error amplification while retaining the regularization yields a slight but measurable improvement (75.98% vs.
 75.91%).
-Finally, by avoiding error-enhancement entirely and preserving the helpful regularization, MaxSup achieves 76.12% accuracy—surpassing LS.
-This result underscores that MaxSup directly addresses LS’s primary shortcoming by consistently applying the intended regularization even when the model’s top-1 prediction is incorrect.
+By fully removing error amplification and faithfully preserving the helpful aspects of LS, our MaxSup achieves 76.12% accuracy, clearly and consistently outperforming LS.
+This result underscores that MaxSup directly tackles LS’s fundamental shortcoming by maintaining a consistent and meaningful regularization signal—even when the top-1 prediction is incorrect.
 3.2 Max Suppression Regularization
-Label Smoothing (LS) suffers from two main limitations: inconsistent regularization and error amplification.
-As discussed in Section 3.1 and illustrated in Table 1, LS penalizes the ground-truth logit $z_{gt}$ even for misclassified examples, thereby unnecessarily widening the gap between $z_{gt}$ and the incorrect top-1 logit.
-To address these critical shortcomings, we propose Max Suppression Regularization (MaxSup), which explicitly penalizes the largest logit $z_{\textit{max}}$ rather than $z_{gt}$.
-This crucial shift ensures uniform regularization across both correct and misclassified samples, effectively eliminating the error-amplification issue seen in LS (Table 1), and preserving the integrity of the ground-truth logit for more stable and robust learning performance.
+Building on our analysis in Section 3.1, we find that Label Smoothing (LS) not only impacts correctly classified samples but also influences misclassifications in unintended and harmful ways.
+Specifically, LS suffers from two main limitations: inconsistent regularization and error amplification.
+As illustrated in Table 1, LS penalizes the ground-truth logit $z_{gt}$ even in misclassified examples, needlessly widening the gap between $z_{gt}$ and the erroneous top-1 logit.
+To resolve these critical shortcomings, we propose Max Suppression Regularization (MaxSup), which explicitly penalizes the largest logit $z_{\textit{max}}$ rather than $z_{gt}$.
+This key design choice ensures uniform regularization across both correct and misclassified samples, effectively eliminating the error-amplification issue in LS (Table 1) and preserving the ground-truth logit’s integrity for more stable, robust learning.
 Definition 3.5.
@@
 By shifting the penalty from $z_{gt}$ (the ground-truth logit) to $z_{\textit{max}}$ (the highest logit), MaxSup avoids unduly suppressing $z_{gt}$ when the model misclassifies, thus overcoming Label Smoothing’s principal shortcoming.
-Logit-Level Formulation of MaxSup
-Building on the logit-level perspective introduced for LS in Section 3.1, we can express $L_{\textit{MaxSup}}$ as:
+Logit-Level Formulation of MaxSup. Building on the logit-level perspective introduced for LS in Section 3.1, we can express $L_{\textit{MaxSup}}$ as:
 $L_{\textit{MaxSup}}\;=\;\alpha\Bigl{(}z_{\textit{max}}\;-\;\tfrac{1}{K}\sum_{k=1}^{K}z_{k}\Bigr{)},$ (8)
@@
 As shown in Table 1, this approach eliminates LS’s error-amplification issue while preserving the intended overconfidence suppression.
-Comparison with Label Smoothing
+Comparison with Label Smoothing.
 MaxSup fundamentally differs from LS in handling correct and incorrect predictions.
 When $z_{gt}=z_{\textit{max}}$, both LS and MaxSup similarly reduce overconfidence.
-However, when $z_{gt}\neq z_{\textit{max}}$, LS continues to shrink $z_{gt}$, widening the gap with the incorrect logit, whereas MaxSup penalizes $z_{\textit{max}}$, preserving $z_{gt}$ from undue suppression.
-As illustrated in Figure 3, this allows the model to recover from mistakes more effectively and avoid reinforcing incorrect predictions.
-Gradient Analysis
-To understand MaxSup’s optimization dynamics, we compute its gradients with respect to each logit $z_{k}$. Specifically,
+However, when $z_{gt}\neq z_{\textit{max}}$, LS shrinks $z_{gt}$, widening the gap with the incorrect logit, whereas MaxSup penalizes $z_{\textit{max}}$, preserving $z_{gt}$ from undue suppression.
+As illustrated in Figure 2, this helps the model recover from mistakes more effectively and avoid reinforcing incorrect predictions.
+Gradient Analysis. To understand MaxSup’s optimization dynamics, we compute its gradients with respect to each logit $z_{k}$. Specifically,
 $\frac{\partial L_{\textit{MaxSup}}}{\partial z_{k}}\;=\;\begin{cases}\alpha\Bigl{(}1-\tfrac{1}{K}\Bigr{)},&\text{if }k=\arg\max(\mathbf{q}),\\ -\tfrac{\alpha}{K},&\text{otherwise}.\end{cases}$ (9)
-Thus, the top-1 logit $z_{\textit{max}}$ is reduced by $\alpha\bigl{(}1-\tfrac{1}{K}\bigr{)}$, while all other logits increase slightly by $\tfrac{\alpha}{K}$.
-In misclassified cases, the ground-truth logit $z_{gt}$ is therefore spared from penalization, thereby avoiding the error-amplification issue seen in LS.
-For completeness, Appendix A provides a full derivation of these gradients, and Figure 2 compares the resulting logit distributions under different regularizers.
-Behavior Across Different Samples
-MaxSup applies a dynamic penalty that depends on the model’s current predictions.
-For high-confidence, correctly classified examples, it behaves similarly to LS by reducing overconfidence, thus effectively mitigating overfitting.
-In contrast, for misclassified or uncertain samples, MaxSup specifically suppresses the incorrect top-1 logit, further safeguarding the ground-truth logit $z_{gt}$.
-This selective strategy preserves an accurate representation of the true class while actively discouraging the propagation of errors.
-As shown in Section 5.1 and Table 5, this promotes more robust decision boundaries and ultimately leads to stronger generalization performance.
-Theoretical Insights and Practical Benefits
-MaxSup provides both theoretical and practical advantages compared to LS.
-Whereas LS applies a uniform penalty to the ground-truth logit regardless of correctness, MaxSup focuses on penalizing only the most confident logit $z_{\textit{max}}$.
-This dynamic adjustment prevents error accumulation in misclassifications, thereby ensuring more stable convergence.
-As a result, MaxSup achieves stronger generalization, exhibits greater robustness to label noise, and performs well on challenging datasets.
-Moreover, as shown in Section 4, MaxSup preserves higher intra-class diversity, which substantially improves transfer learning performance (Table 3) and yields more interpretable activation maps (Figure 3).
-4 Analysis of MaxSup’s Learning Benefits
-MaxSup simultaneously promotes inter-class separability and intra-class variation, both essential for robust classification and effective feature transfer.
-In this section, we explore how MaxSup achieves these objectives and contrast its effectiveness with alternative regularization methods.
-4.1 Intra-Class Variation and Transferability
-As noted in Section 3.1, Label Smoothing (LS) primarily restricts overconfidence when the ground-truth class is correctly predicted, inadvertently causing error enhancement for misclassified samples.
-This selective penalty can overly compress intra-class diversity.
-In contrast, MaxSup uniformly penalizes the top-1 logit in both correct and incorrect cases, eliminating LS’s error-enhancement component and thus preserving more fine-grained distinctions within each class.
-Table 2 compares intra-class variation $\bar{d}_{\text{within}}$ and inter-class separability $R^{2}$ (Kornblith et al., 2021) for a ResNet-50 model trained on ImageNet-1K.
-Although all regularization strategies reduce $\bar{d}_{\text{within}}$ relative to the baseline, MaxSup shows the smallest reduction, implying stronger retention of within-class variability—often correlated with improved generalization and transferability.
-The benefits of this richer intra-class structure appear clearly in Table 3, where linear transfer performance on CIFAR-10 is reported.
-Although LS and Logit Penalty improve ImageNet accuracy, they diminish transfer accuracy by over-suppressing informative features.
-In contrast, MaxSup preserves transfer performance near that of the baseline, suggesting it retains crucial, discriminative features that generalize effectively to downstream tasks.
-Table 2: Metrics for feature representation quality using ResNet-50 trained on ImageNet-1K.
-We report intra-class variation ($\bar{d}_{\text{within}}$) and inter-class separability ($R^{2}$), both of which benefit from higher values.
-Although all methods reduce $\bar{d}_{\text{within}}$ relative to the baseline, MaxSup preserves the most within-class diversity.
+Thus, the top-1 logit $z_{\textit{max}}$ is reduced by $\alpha\bigl{(}1-\tfrac{1}{K}\bigr{)}$, while all other logits slightly increase by $\tfrac{\alpha}{K}$.
+In misclassified cases, the ground-truth logit $z_{gt}$ is spared from penalization, avoiding the error-amplification issue seen in LS.
+For completeness, Appendix A provides the full gradient derivation.
+Behavior Across Different Samples.
+MaxSup applies a dynamic penalty based on the model’s current predictions.
+For high-confidence, correctly classified examples, it behaves similarly to LS by reducing overconfidence, effectively mitigating overfitting.
+In contrast, for misclassified or uncertain samples, MaxSup aggressively suppresses the incorrect top-1 logit, further safeguarding the ground-truth logit $z_{gt}$.
+This selective strategy preserves a faithful and reliable representation of the true class while actively discouraging error propagation.
+As shown in Section 4.2 and Table 5, this promotes more robust decision boundaries and leads to stronger generalization.
+Theoretical Insights and Practical Benefits.
+MaxSup provides both theoretical and practical advantages over LS.
+Whereas LS applies a uniform penalty to the ground-truth logit regardless of correctness, MaxSup penalizes only the most confident logit $z_{\textit{max}}$.
+This dynamic adjustment robustly prevents error accumulation in misclassifications, ensuring more stable convergence.
+As a result, MaxSup generalizes better and achieves strong performance on challenging datasets.
+Moreover, as shown in Section 4.1, MaxSup preserves greater intra-class diversity, substantially improving transfer learning (Table 3) and yielding more interpretable activation maps (Figure 2).
+4 Experiments
+We begin by examining how MaxSup improves feature representations, then evaluate it on large-scale image classification and semantic segmentation tasks.
+Finally, we visualize class activation maps to illustrate the practical benefits of MaxSup.
+4.1 Analysis of MaxSup’s Learning Benefits
+Having established how MaxSup addresses Label Smoothing’s (LS) principal shortcomings (Section 3.1), we now demonstrate its impact on inter-class separability and intra-class variation—two properties essential for accurate classification and effective transfer learning.
+Table 2: Feature quality metrics for ResNet-50 on ImageNet-1K.
+While all methods reduce $\bar{d}_{\text{within}}$ (intra-class variation) relative to the baseline, MaxSup retains the greatest within-class diversity.
+Both $\bar{d}_{\text{within}}$ and $R^{2}$ benefit from higher values.
 Method $\bar{d}_{\text{within}}\!\uparrow$ $R^{2}\!\uparrow$
 Train Val Train Val
-Baseline 0.3114 0.3313 0.4025 0.4451
-LS 0.2632 0.2543 0.4690 0.4611
-OLS 0.2707 0.2820 0.5943 0.5708
-Zipf’s 0.2611 0.2932 0.5522 0.4790
-MaxSup 0.2926 0.2998 0.5188 0.4972
-Logit Penalty 0.2840 0.3144 0.6448 0.6024
-4.2 Impact of Logit Regularization
-Different regularization methods impose distinct constraints on the logit space, thereby shaping the model’s representational capacity (Kornblith et al., 2021).
-Among these approaches, Logit Penalty and MaxSup both act directly on logits but differ fundamentally in how they apply regularization.
-Logit Penalty operates by minimizing the $\ell_{2}$-norm of the entire logit vector, causing a global reduction in logit magnitudes that often induces sparsity.
-This uniform shrinkage can limit intra-class variation, thereby weakening the model’s ability to transfer features to downstream tasks.
-In contrast, MaxSup targets only the largest (top-1) logit, nudging it closer to the average logit.
-By selectively penalizing only the most confident prediction, MaxSup avoids universal shrinkage and preserves richer intra-class diversity, a property crucial for effective transferability.
-Figure 2 illustrates the distribution of logits under various regularizers.
-Logit Penalty yields a narrower logit range, reflecting excessive sparsity and aligning with its lower transfer performance (Table 3).
-By comparison, MaxSup maintains broader logit distributions, thereby retaining the fine-grained feature distinctions needed to excel on downstream tasks.
-Table 3: Validation performance on CIFAR-10 with a linear probe using $l_{2}$-regularized multinomial logistic regression.
-Although Label Smoothing and Logit Penalty improve ImageNet accuracy, they substantially degrade transfer accuracy compared to MaxSup.
-Method Linear Transfer Acc.
-Baseline 0.8143
-Label Smoothing 0.7458
-Logit Penalty (Dauphin & Cubuk, 2021) 0.7242
-MaxSup 0.8102
-Figure 2: Logit density plots under three different regularization strategies: MaxSup, Logit Penalty, and standard Cross Entropy.
-Logit Penalty induces a narrower logit distribution, reflecting excessive shrinkage that reduces intra-class variation.
-By contrast, MaxSup preserves a broader range of logits and thus richer representations.
-5 Experiments
+Baseline 0.311 0.331 0.403 0.445
+LS 0.263 0.254 0.469 0.461
+OLS 0.271 0.282 0.594 0.571
+Zipf-LS 0.261 0.293 0.552 0.479
+MaxSup 0.293 0.300 0.519 0.497
+Logit Penalty 0.284 0.314 0.645 0.602
+4.1.1 Intra-Class Variation and Transferability
+As noted in Section 3.1, Label Smoothing (LS) primarily curbs overconfidence for correctly classified samples but inadvertently triggers error amplification in misclassifications.
+This uneven penalization can overly compress intra-class feature representations.
+By contrast, MaxSup uniformly penalizes the top-1 logit, whether the prediction is correct or incorrect, thereby eliminating LS’s error-amplification effect and preserving finer distinctions within each class.
+Table 2 compares intra-class variation ($\bar{d}_{\text{within}}$) and inter-class separability ($R^{2}$[13] for ResNet-50 trained on ImageNet-1K.
+Although all investigated regularizers decrease $\bar{d}_{\text{within}}$ relative to a baseline, MaxSup yields the smallest reduction, indicating a stronger retention of subtle within-class diversity—widely associated with enhanced generalization and improved transfer performance.
+These benefits are further underscored by the linear-probe transfer accuracy on CIFAR-10 (Table 3).
+While LS and Logit Penalty each boost ImageNet accuracy, both degrade transfer accuracy, likely by suppressing informative and transferable features.
+By contrast, MaxSup preserves near-baseline performance, implying that it maintains rich discriminative information crucial for downstream tasks.
+For extended evaluations on diverse datasets, see Table 9 in the appendix.
+Table 3: Linear-probe transfer accuracy on CIFAR-10 (higher is better).
+Method Acc.
+Baseline 0.814
+Label Smoothing 0.746
+Logit Penalty 0.724
+MaxSup 0.810
+4.1.2 Connection to Logit Penalty
+As detailed in Section 3, both Label Smoothing (LS) variants and MaxSup impose penalties directly at the logit level, aligning with the perspective that various regularizers influence a model’s representational capacity via distinct logit constraints [13].
+Within this family of techniques, Logit Penalty and MaxSup both address the maximum logit, yet diverge fundamentally in their specific methods of regularization.
+Logit Penalty minimizes the $\ell_{2}$-norm of the entire logit vector, inducing a global contraction that can improve class separation but also reduce intra-class diversity, potentially hindering downstream transfer.
+By contrast, MaxSup focuses exclusively on the top-1 logit, gently nudging it closer to the mean logit.
+Because only the highest-confidence prediction is penalized, MaxSup avoids the uniform shrinkage observed in Logit Penalty, preserving richer intra-class variation—a property essential for robust transfer.
+Further insights into this behavior can be found in Appendix N, where logit-value histograms illustrate how each method affects the logit distribution.
 4.2 Evaluation on ImageNet Classification
-In this section, we assess the effectiveness of MaxSup on ImageNet-1K, comparing its performance against standard Label Smoothing and related variants.
+Next, we compare MaxSup to standard Label Smoothing (LS) and various LS extensions on the large-scale ImageNet-1K dataset.
 4.2.1 Experiment Setup
-Model Training Configurations
-We conduct extensive experiments with both CNN and Transformer models, including the ResNet family (He et al., 2016), MobileNetV2 (Sandler et al., 2018), and DeiT-Small (Touvron et al., 2021), all thoroughly evaluated on the large-scale ImageNet dataset (Krizhevsky et al., 2012) for comprehensive performance analysis.
-For ResNet Series models, we train for 200 epochs using stochastic gradient descent (SGD) with momentum 0.9, a weight decay of $1\times 10^{-4}$, and a batch size of 2048.
-The initial learning rate is set to 0.85 and scheduled via cosine annealing.
-[^2] Additional training hyperparameters follow the FFCV training scripts in https://github.com/libffcv/ffcv.
-See Appendix E for further details on training setups.
-We also evaluate ResNet-based CNNs on CIFAR-100.
-Here, we use an initial learning rate of 0.1, reducing it by a factor of 5 at the 60th, 120th, and 160th epochs.
-We train for 200 epochs with a batch size of 128, weight decay of $5\times 10^{-4}$, and Nesterov momentum set to 0.9.
-For DeiT-Small, we employ the official implementation and train from scratch without knowledge distillation.
-Although the original DeiT paper emphasizes distillation, we exclude it to provide a clearer, unbiased assessment of MaxSup’s contributions.
-We also omit CutMix and Mixup to retain the same optimization objective.
-Hyperparameters for Compared Methods
-We compare Max Suppression Regularization against multiple Label Smoothing variants, including Zipf Label Smoothing (Liang et al., 2022) and Online Label Smoothing (Zhang et al., 2021).
-When official implementations are available, we use them directly; otherwise, we follow the respective papers’ descriptions closely to ensure fair comparisons.
-All training hyperparameters are kept identical to those of the baseline models, except for algorithm-specific settings and necessary adjustments.
-Additionally, we employ a linearly increasing $\alpha$ scheduler, which generally benefits training and stability; see Appendix F for details.
-This scheduler is applied to both MaxSup and standard Label Smoothing by default to maintain consistency.
+Model Training Configurations.We evaluate both convolutional (ResNet[11], MobileNetV2 [25]) and transformer (DeiT-Small [33]) architectures on ImageNet [15].
+For the ResNet Series, we train for 200 epochs using stochastic gradient descent (SGD) with momentum0.9, weight decay of $1\times 10^{-4}$, and a batch size of 2048.
+The initial learning rate is 0.85 and is annealed via a cosine schedule.
+[^2] Additional training hyperparameters follow the FFCV scripts at https://github.com/libffcv/ffcv.
+See Appendix E for further details.
+We also test ResNet variants on CIFAR-100 with a conventional setup: an initial learning rate of 0.1 (reduced fivefold at epochs 60, 120, and 160), training for 200 epochs with batch size 128 and weight decay $5\times 10^{-4}$.
+For DeiT-Small, we use the official codebase [33], training from scratch without knowledge distillation to isolate MaxSup’s contribution.
+CutMix and Mixup are disabled to ensure the model optimization objective remains unchanged.
+Hyperparameters for Compared Methods.We compare Max Suppression Regularization against a range of LS extensions, including Zipf Label Smoothing[19] and Online Label Smoothing [41].
+Where official implementations exist, we adopt them directly; otherwise, we follow the methodological details provided in each respective paper.
+Except for any method-specific hyperparameters, all other core training settings remain identical to the baselines.
+Furthermore, both MaxSup and standard LS employ a linearly increasing $\alpha$-scheduler for improved training stability (see Appendix F).
+This ensures a fair comparison under consistent and reproducible training protocols.
 4.2.2 Experiment Results
-ConvNet Comparison
-Table 4 summarizes the performance of MaxSup alongside various smoothing and self-distillation methods on both ImageNet and CIFAR-100.
-Across all tested convolutional architectures, MaxSup achieves the highest accuracy among label smoothing–based regularizers.
-By contrast, OLS (Zhang et al., 2021) and Zipf-LS (Liang et al., 2022) yield less consistent gains, suggesting their reported empirical benefits may depend heavily on specific training schedules.
-In our reproductions of OLS and Zipf-LS, we follow the authors’ original codebases and method-specific hyperparameters but do not adopt their complete training recipes.
-For example, the OLS paper uses a step learning-rate scheduler over 250 epochs with an initial rate of 0.1, while Zipf-LS trains for 100 epochs under a separate set of hyperparameters.
-Our results underscore the robustness of MaxSup across different training setups, in contrast to the more scheme-dependent improvements noted for OLS and Zipf-LS.
-Table 4: Comparison of classic convolutional neural networks on ImageNet and CIFAR-100.
-Results are reported as “mean ± std” (percentage).
-Bold entries highlight the best performance; underlined entries mark the second best.
-(Methods with denote code adaptations from official repositories; see text for details.)
+ConvNet Comparison.
+Table4 shows results for MaxSup alongside various label-smoothing and self-distillation methods on both ImageNet and CIFAR-100 benchmarks.
+Across all convolutional architectures tested, MaxSup consistently delivers the highest top-1 accuracy among label-smoothing approaches.
+By contrast, OLS [41] and Zipf-LS [19] exhibit less stable gains, suggesting their effectiveness may heavily hinge on specific training protocols.
+To reproduce OLS and Zipf-LS, we apply the authors’ official codebases and hyperparameters but do not replicate their complete training recipes (e.g., OLS trains for 250 epochs with a step-scheduled learning rate of 0.1, and Zipf-LS uses 100 epochs with distinct hyperparameters).
+Even under these modified settings, MaxSup remains robust, highlighting its effectiveness across a variety of training schedules—unlike the more schedule-sensitive improvements noted for OLS and Zipf-LS.
+Table 4: Performance comparison of classical convolutional networks on ImageNet and CIFAR-100.
+All results are shown as “mean ± std” (percentage).
+Bold highlights the best performance; underlined marks the second best.
+(Methods labeled with indicate code adapted from official repositories; see the text for additional details.)
 Method ImageNet CIFAR-100
@@
 Logit Penalty 68.48±0.10 76.73±0.10 77.20±0.15 71.13±0.10 76.41±0.15 78.90±0.16 78.89±0.21 69.46±0.08
-DeiT Comparison
-Table 5 compares various regularization techniques for DeiT-Small on ImageNet.
-MaxSup achieves an accuracy of $76.49\%$, surpassing Label Smoothing by $0.41$ percentage points.
-Label Smoothing variants such as Zipf’s and OLS yield only marginally higher or comparable performance relative to standard LS, suggesting that these approaches may be less effective for vision transformer architectures—potentially due to their reliance on extensive data augmentation schemes.
-In contrast, MaxSup consistently outperforms both standard LS and its variants, indicating its stronger ability to enhance feature representations without additional data manipulations.
-These findings underscore MaxSup’s robustness across distinct model architectures, especially in settings where other regularization methods show limited effectiveness.
-Table 5: Accuracy (%) comparison on DeiT-Small (Touvron et al., 2021) using different Label Smoothing variants.
-Results are reported as “mean ± std”; parentheses indicate absolute improvement over the baseline.
-Method Accuracy (Mean) Std
+Table 5: DeiT-Small top-1 accuracy (%), reported as mean $\pm$ standard deviation. Values in parentheses indicate absolute improvements over the baseline.
+Method Mean Std
 Baseline 74.39 0.19
@@
 MaxSup 76.49 ($+2.10$) 0.12
-Table 6: Semantic segmentation results on the ADE20K validation set.
-Models are pretrained on ImageNet-1K and then fine-tuned with UperNet (Xiao et al., 2018) for enhanced performance.
-We report mean Intersection over Union (mIoU) under multi-scale (MS) testing for comprehensive evaluation.
-Backbone Method mIoU (MS)
-DeiT-Small (Touvron et al., 2021) Baseline 42.1
+DeiT Comparison.
+Table 5 summarizes performance for DeiT-Small on ImageNet across various regularization strategies.
+Notably, MaxSup attains a top-1 accuracy of 76.49%, surpassing standard Label Smoothing by 0.41%.
+In contrast, LS variants such as Zipf-LS and OLS offer only minor gains over LS, implying that their heavy reliance on data augmentation may limit their applicability to vision transformers.
+By outperforming both LS and its variants without additional data manipulations, MaxSup demonstrates robust feature enhancement.
+These findings underscore MaxSup’s adaptability to different architectures and emphasize its utility in scenarios where conventional label-smoothing methods yield limited benefits.
+Fine-Grained Classification.
+Beyond ImageNet, we evaluate MaxSup on demanding fine-grained benchmarks such as CUB-200 [34] and Stanford Cars [14].
+As shown in Table 11, MaxSup outperforms other LS-based methods by more effectively capturing subtle within-class variations, which are critical for success in fine-grained recognition scenarios.
+Ablation on the Weight Schedule.
+We also systematically investigate how different $\alpha$ scheduling strategies impact MaxSup’s performance.
+Empirical results indicate that MaxSup consistently maintains high accuracy across a wide range of schedules, further underscoring its robustness against hyperparameter changes.
+For additional details and discussions, refer to Appendix F.
+Table 6: Semantic segmentation (multi-scale) on ADE20K using UperNet. All models are pretrained on ImageNet-1K; mIoU reported as percentage.
+Backbone Method mIoU
+DeiT-Small Baseline 42.1
 Label Smoothing 42.4 ($+0.3$)
 MaxSup 42.8 ($+0.7$)
-Figure 3: Class activation maps generated by Grad-CAM (Selvaraju et al., 2019) for DeiT-Small models trained with MaxSup (2nd row), Label Smoothing (3rd row), and a standard Baseline (4th row).
+4.3 Evaluation on Semantic Segmentation
+We further investigate MaxSup’s applicability to downstream tasks by evaluating its performance on semantic segmentation using the widely adopted MMSegmentation framework.
+[^3] https://github.com/open-mmlab/mmsegmentation Specifically, we adopt the UperNet [36] architecture with a DeiT-Small backbone, trained on ADE20K.
+Models pretrained on ImageNet-1K with either MaxSup or Label Smoothing are then fine-tuned under the same cross-entropy objective (Section 4.2.2).
+Table 6 shows that initializing with MaxSup-pretrained weights yields an mIoU of $42.8\%$, surpassing the $42.4\%$ achieved by Label Smoothing.
+This improvement indicates that MaxSup fosters more discriminative feature representations conducive to dense prediction tasks.
+By more effectively capturing class boundaries and within-class variability, MaxSup promotes stronger segmentation results, underscoring its potential to deliver features that are both transferable and highly robust.
+Figure 2: Grad-CAM [27] visualizations for DeiT-Small models under three training setups: MaxSup (2nd row), Label Smoothing (3rd row), and a baseline (4th row).
 The first row shows the original images.
-Compared with Label Smoothing, MaxSup more effectively suppresses distractions from non-target objects and preserves key features of the target class, thereby reducing instances in which the model partially or completely focuses on irrelevant regions.
-5.2 Evaluation on Semantic Segmentation
-To further assess the transferability of MaxSup to downstream tasks, we evaluate its performance on semantic segmentation using the MMSegmentation framework.
-[^3] https://github.com/open-mmlab/mmsegmentation Specifically, we employ the UperNet architecture (Xiao et al., 2018) with a DeiT-Small backbone, trained on ADE20K.
-We compare backbones trained with Label Smoothing and MaxSup (on ImageNet-1K) against a baseline, following the same setup as in Section 5.1.2.
-During fine-tuning, all models use a standard cross-entropy loss.
-As shown in Table 6, MaxSup achieves a mean Intersection over Union (mIoU) of $42.8\%$, surpassing the $42.4\%$ obtained with Label Smoothing.
-These findings further highlight the improved feature representations afforded by MaxSup in downstream tasks such as semantic segmentation.
+Compared to Label Smoothing, MaxSup more effectively filters out non-target regions and highlights essential features of the target class, reducing instances where the model partially or entirely focuses on irrelevant areas.
 4.4 Visualization via Class Activation Maps
-To assess how MaxSup influences model decision-making relative to Label Smoothing (LS), we employ Gradient-weighted Class Activation Mapping (Grad-CAM) (Selvaraju et al., 2019).
-Grad-CAM produces class-discriminative localization maps that highlight the regions most relevant to each classification decision.
-We conduct comprehensive experiments on the DeiT-Small model under three distinct training setups: MaxSup (second row), Label Smoothing (third row), and standard Cross-Entropy (CE) as a baseline (fourth row).
-As illustrated in Figure 3, MaxSup-trained models exhibit a distinct advantage in effectively mitigating distractions caused by non-target salient objects in the background (e.g., a pole in the ‘Bird’ image, a tube in the ‘Goldfish’ image, and a cap in the ‘House Finch’ image).
-By contrast, LS-trained models often lose focus or erroneously attend to background objects, further reflecting the detrimental influence of LS’s Error-Enhancement term and its impact on feature learning (Please see Figures 3).
-Moreover, MaxSup preserves a wider range of relevant object features, as evidenced in the ‘Shark’ and ‘Monkey’ examples, where LS-trained models fail to capture key details (fins and tails).
-These observations align with the analysis in Appendix G, underscoring that MaxSup better retains rich intra-class information.
-Consequently, MaxSup-trained models yield more accurate and reliable classifications by leveraging these detailed feature representations.
-5.4 Quantitative Analysis of CAM Overlays
-To further quantify the benefits of MaxSup, we measure precision and recall for Grad-CAM overlays on target object regions.
-Our results indicate that MaxSup consistently attains higher localization precision while maintaining or improving recall, leading to more accurate and comprehensive feature activation.
-Qualitative assessments similarly reveal that MaxSup-trained models exhibit both a more focused and more extensive attention map, resulting in more reliable and interpretable predictions.
-These findings demonstrate that MaxSup not only enhances feature representation but also contributes to more trustworthy AI systems.
+To better understand how MaxSup fundamentally differs from Label Smoothing (LS) in guiding model decisions, we employ Gradient-weighted Class Activation Mapping (Grad-CAM) [27], which highlights regions most influential for each prediction.
+We evaluate DeiT-Small under three training setups: MaxSup (second row), LS (third row), and a baseline with standard cross-entropy (fourth row).
+As illustrated in Figure 2, MaxSup-trained models more effectively suppress background distractions than LS, which often fixates on unrelated objects—such as poles in “Bird,” tubes in “Goldfish,” and caps in “House Finch.” This behavior reflects LS’s error-enhancement mechanism, which can misdirect attention.
+Moreover, MaxSup retains a wider spectrum of salient features, as exemplified in the Shark” and Monkey” images, where LS-trained models often omit crucial semantic details (e.g., fins, tails, or facial contours).
+These findings align with our analysis in Appendix J, clearly demonstrating that MaxSup preserves richer intra-class information.
+Consequently, MaxSup-trained models produce more accurate and consistent predictions by effectively leveraging fine-grained object cues.
+Further quantitative Grad-CAM overlay metrics (e.g., precision and recall for target regions) confirm that MaxSup yields more focused and comprehensive activation maps, further underscoring its overall efficacy.
 5 Conclusion
-In this work, we carefully examined the root causes of Label Smoothing’s (LS) shortcomings and introduced Max Suppression Regularization (MaxSup) as a targeted remedy.
-Our analysis shows that LS can unintentionally promote overconfidence in misclassified samples by applying insufficient regularization to erroneous top-1 logits.
-In contrast, MaxSup effectively addresses this by consistently penalizing the most confident logit, regardless of prediction correctness.
-Through extensive experiments and in-depth analyses, we demonstrate that MaxSup not only improves accuracy but also preserves richer intra-class variation and significantly enhances inter-class separability.
-Consequently, models trained with MaxSup capture finer-grained information about individual samples, ultimately leading to stronger transfer learning capabilities.
-Class activation maps further reveal that MaxSup directs model attention more accurately toward salient parts of target objects, effectively mitigating distractions from background elements.
-Limitations & Future Work
-Although our findings validate MaxSup’s effectiveness, several directions merit further investigation.
-Prior work (Müller et al., 2019) shows that teachers trained with LS can degrade performance in Knowledge Distillation (Hinton, 2015), and Guo et al.
-(2024) suggests LS accelerates convergence via conditioning number analysis.
-Future research could explore MaxSup’s impact on Knowledge Distillation workflows and its influence on training convergence.
-Additionally, recent studies (Súkeník et al., 2024; Garrod & Keating, 2024) indicate that $\ell_{2}$ regularization biases final layer features and weights toward lower-rank solutions than those typically associated with neural collapse.
-Investigating how MaxSup interacts with these low-rank biases and whether it leads to similarly optimal or novel solutions is another intriguing avenue for future work.
-Impact Statement
-This paper advances machine learning regularization techniques through an in-depth analysis of Label Smoothing and a proposed alternative, Max Suppression Regularization.
-By reinforcing generalization and transferability on widely used computer vision tasks, MaxSup has the potential to benefit a broad array of real-world applications, improving robustness without introducing known negative consequences.
-Thus, we anticipate a net positive impact, offering practitioners a clearer understanding of network regularization and access to more effective training methodologies.
-Acknowledgements
-This research was supported by the University of Washington Graduate School’s GSFEI Top Scholar Award and the Start-Up Funding awarded to Prof.
-Zhi-Qi Cheng.
-Additional support was provided by ELSA – European Lighthouse on Secure and Safe AI (No.
-101070617).
+We examined the shortcomings of Label Smoothing (LS) and introduced Max Suppression Regularization (MaxSup) as a targeted and practical remedy.
+Our analysis shows that LS can unintentionally heighten overconfidence in misclassified samples by failing to sufficiently penalize incorrect top-1 logits.
+In contrast, MaxSup uniformly penalizes the highest logit, regardless of prediction correctness, thereby effectively eliminating LS’s error amplification.
+Extensive experiments demonstrate that MaxSup not only improves accuracy but also preserves richer intra-class variation and enforces sharper inter-class boundaries, leading to more nuanced and transferable feature representations and superior transfer performance.
+Moreover, class activation maps confirm that MaxSup better attends to salient object regions, reducing focus on irrelevant background elements.
+Limitations.
+Although promising, several questions remain.
+Prior work[21] notes that LS-trained teachers may degrade knowledge distillation [12], and Guo et al.
+[9] suggests LS accelerates convergence via conditioning.
+Examining MaxSup’s role in distillation and its impact on training dynamics would clarify these effects.
+Recent studies [31, 7] also show that $\ell_{2}$ regularization biases final-layer features toward low-rank solutions, raising questions about whether MaxSup behaves similarly.
+Impact.
+In practical applications, MaxSup shows strong promise for systems demanding robust generalization and efficient transfer, and we have not observed any additional adverse effects or trade-offs.
+By offering researchers and practitioners both a clearer understanding of LS’s limitations and a straightforward, computationally light, and easily integrable method to overcome them, MaxSup may help guide the development of more reliable and interpretable deep learning models.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4.
+Justification: The paper’s abstract and introduction outline the main contributions—including the identification of Label Smoothing (LS) shortcomings and the proposal of Max Suppression (MaxSup)—and these claims align with the theoretical and experimental sections.
+  5. 2.
+    Limitations
+  6. Question: Does the paper discuss the limitations of the work performed by the authors?
+  7. Answer: [Yes]
+  8.
+Justification: A dedicated “Limitations” portion (or equivalent discussion) is provided, acknowledging possible extensions (e.g., knowledge distillation scenarios) and other open questions (e.g., interactions with $\ell_{2}$ regularization).
+  9. 3.
+    Theory assumptions and proofs
+  10. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  11. Answer: [Yes]
+  12.
+Justification: The paper includes formal statements and proofs in the main text and/or appendix (Lemma/Theorem with proofs in the supplementary material).
+All assumptions are clearly stated, and references are provided.
+  13. 4.
+    Experimental result reproducibility
+  14. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results?
+  15. Answer: [Yes]
+  16. Justification: The main text and appendix provide training pipelines, hyperparameters, datasets, and references to the code. Full details (batch sizes, learning rates, etc.) are included.
+  17. 5.
+    Open access to data and code
+  18. Question: Does the paper provide open access to the data and code?
+  19. Answer: [Yes]
+  20. Justification: The code is released (anonymized if needed), and the datasets used (ImageNet, CIFAR, etc.) are publicly available under their respective standard licenses.
+  21. 6.
+    Experimental setting/details
+  22. Question: Does the paper specify all the training and test details necessary to understand the results?
+  23. Answer: [Yes]
+  24. Justification: Section 4.2.1 and the appendix detail the setup (optimizers, data splits, learning rates, etc.). The authors specify how they selected key hyperparameters.
+  25. 7.
+    Experiment statistical significance
+  26. Question: Does the paper report error bars or statistical significance for the experiments?
+  27. Answer: [Yes]
+  28. Justification: Tables report “mean $\pm$ std” from multiple runs, reflecting the variability due to initialization or training seeds. This is shown in all main experimental tables.
+  29. 8.
+    Experiments compute resources
+  30. Question: Does the paper provide sufficient information about compute resources?
+  31. Answer: [Yes]
+  32.
+Justification: The text or appendix indicates GPU usage (e.g., ResNet on cluster GPUs), approximate training duration, and other relevant details.
+Though high-level, it suffices to gauge feasibility.
+  33. 9.
+    Code of ethics
+  34. Question: Does the research conform with the NeurIPS Code of Ethics?
+  35. Answer: [Yes]
+  36. Justification: The work adheres to standard academic norms, uses publicly available datasets, and presents no known ethical concerns or conflicts with the NeurIPS Code of Ethics.
+  37. 10.
+    Broader impacts
+  38. Question: Does the paper discuss both potential positive and negative societal impacts?
+  39. Answer: [Yes]
+  40.
+Justification: The “Impact” statement addresses potential benefits (improved accuracy and transfer, leading to more robust systems) and acknowledges that misuses are minimal given the method’s purely algorithmic nature.
+  41. 11.
+    Safeguards
+  42. Question: Does the paper describe safeguards for high-risk data or models?
+  43. Answer: [N/A]
+  44.
+Justification: The paper does not involve high-risk data (e.g., private user info) or high-risk models (e.g., generative LLMs).
+Standard ImageNet/CIFAR usage and training code are of no particular misuse risk.
+  45. 12.
+    Licenses for existing assets
+  46. Question: Are the creators of assets properly credited, and the licenses mentioned?
+  47. Answer: [Yes]
+  48. Justification: The paper cites and credits publicly available code or datasets (ImageNet, CIFAR, etc.) with references to their original licenses or terms of service.
+  49. 13.
+    New assets
+  50. Question: Are newly introduced assets well documented?
+  51. Answer: [N/A]
+  52.
+Justification: No new data or special code libraries are introduced beyond the regular code release.
+The approach modifies existing code for training frameworks but does not constitute a new dataset or model asset.
+  53. 14.
+    Crowdsourcing and research with human subjects
+  54. Question: Are there human subjects or crowdsourcing experiments, with instructions and compensation described?
+  55. Answer: [N/A]
+  56. Justification: The work involves no human subjects or crowdsourcing tasks.
+  57. 15.
+    Institutional review board (IRB) approvals or equivalent
+  58. Question: Does the paper discuss IRB approvals for human-subjects research?
+  59. Answer: [N/A]
+  60. Justification: The paper does not involve human subjects; no IRB is necessary.
+  61. 16.
+    Declaration of LLM usage
+  62. Question: Does the paper describe usage of LLMs if it is essential to core methods in this research?
+  63. Answer: [Yes]
+  64.
+Justification: We used a Large Language Model (LLM) solely for writing and polishing the paper’s text.
+The LLM was not involved in designing, conducting, or analyzing the experiments, nor in developing the core algorithmic contributions.
 Appendix A Proof of Lemma 3.2
@@
 Algorithm 1 Gradient Descent with Max Suppression (MaxSup)
-[…]
+1:Training set $D=\{(\mathbf{x}^{(i)},\mathbf{y}^{(i)})\}_{i=1}^{N}$; learning rate $\eta$; number of iterations $T$; smoothing parameter $\alpha$; a neural network $f_{\theta}(\cdot)$; batch size $B$; total classes $K$.
+2:Initialize network weights $\theta$ (e.g., randomly).
+3:for $t=1$ to $T$ do
+4:    // Each iteration processes mini-batches of size $B$.
+5:    for each mini-batch $\{(\mathbf{x}^{(j)},\mathbf{y}^{(j)})\}_{j=1}^{B}$ in $D$ do
+6:       Compute logits: $\mathbf{z}^{(j)}\leftarrow f_{\theta}(\mathbf{x}^{(j)})$ for each sample in the batch
+7:       Compute predicted probabilities: $\mathbf{q}^{(j)}\leftarrow\text{softmax}(\mathbf{z}^{(j)})$
+8:       Compute cross-entropy loss:
 $L_{\mathrm{CE}}\leftarrow\frac{1}{B}\sum_{j=1}^{B}H\bigl{(}\mathbf{y}^{(j)},\mathbf{q}^{(j)}\bigr{)}$
-   // MaxSup component: penalize the top-1 logitFor each sample $j$:
+9:          // MaxSup component: penalize the top-1 logit
+10:       For each sample $j$:
 $z_{\textit{max}}^{(j)}\;=\;\max_{k\in\{1,\dots,K\}}z_{k}^{(j)},\quad\bar{z}^{(j)}\;=\;\frac{1}{K}\sum_{k=1}^{K}z_{k}^{(j)}$
 $L_{\mathrm{MaxSup}}\leftarrow\frac{1}{B}\sum_{j=1}^{B}\bigl{[}z_{\textit{max}}^{(j)}-\bar{z}^{(j)}\bigr{]}$
-Total loss:
+11:       Total loss:
 $L\;\leftarrow\;L_{\mathrm{CE}}\;+\;\alpha\,L_{\mathrm{MaxSup}}$
-Update parameters:
+12:       Update parameters:
 $\theta\;\leftarrow\;\theta-\eta\,\nabla_{\theta}\,L$
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
+13:    end for
+14:end for
 Appendix D Pseudo Code
@@
 MaxSup $\alpha\Bigl{(}z_{\textit{max}}-\tfrac{1}{K}\sum_{k}z_{k}\Bigr{)}$ 76.12 76.58
-Appendix G Visualization of the Learned Feature Space
+Appendix G Extended Evaluation of Linear Transferability on Different Datasets
+To further demonstrate the substantial improvement in feature representation compared to other methods, we further compare the linear transfer accuracies of different methods on a broader range of datasets in Table 9
+Table 9: Validation performance of different methods, evaluated using multinomial logistic regression with l2 regularization.
+Although Label Smoothing and OLS improve ImageNet accuracy, they substantially degrade transfer accuracy compared to MaxSup.
+Following [13], we selected from 45 logarithmically spaced values between $10^{-6}$ and $10^{5}$.
+Notably, the search range is larger than the search range used in Table 3, thus leading to higher overall accuracies on CIFAR10.
+Datasets CIFAR10 CIFAR100 CUB Flowers Foods Pets
+CE 91.74 75.35 70.21 90.96 72.44 92.30
+LS 90.14 71.28 64.50 84.84 67.76 91.96
+OLS 90.29 73.13 67.86 87.47 69.34 92.21
+MaxSup 91.00 73.93 67.29 88.84 70.94 92.93
+Appendix H Extended Comparison to More Label Smoothing Alternatives
+We have included Confidence Penalty [23] and Adaptive Label Smooothing with Self-Knowledge [18] for an extended comparison in Table 10.
+We follow their recommended hyperparameter settings: weight coefficient is set to 0.1 for confidence penalty, and the checkpoint with the highest validation accuracy checkpoint is treated as teacher for Adaptive Label Smoothing.
+From Table 10, we see that MaxSup outperforms both Confidence Penalty and Adaptive LS on ImageNet with ResNet-50.
+This aligns with our theoretical analysis that selectively penalizing $z_{\mathrm{max}}$ yields a more consistent regularization than penalizing all logits (Confidence Penalty) or dynamically smoothing the label distribution with self knowledge (Adaptive LS).
+Table 10: Comparison of classic convolutional neural networks on ImageNet and CIFAR-100.
+Results are reported as “mean ± std” (percentage).
+Bold entries highlight the best performance; underlined entries mark the second best.
+(Methods with denote code adaptations from official repositories; see text for details.)
+Method ImageNet
+ResNet-50
+Baseline 76.41±0.10
+Label Smoothing 76.91±0.11
+[…]
+$\text{OLS}^{*}$ 77.23±0.21
+MaxSup 77.69±0.07
+Logit Penalty 76.73±0.10
+Confidence Penalty 76.58±0.12
+Adaptive Label Smoothing 77.16±0.15
+Appendix I Classification on Fine-Grained Datasets
+We have further compared MaxSup and other methods on CUB-200 and Stanford Cars for fine-grained classification in Table 11.
+The performance of MaxSup again supports its superiority against existing approaches.
+Table 11: Classification on CUB and Cars Datasets.
+Method CUB[34] Cars[14]
+no_ls 80.88 90.27
+ls 81.96 91.64
+ols 82.33 91.96
+zipfs 81.40 90.99
+max_sup 82.53 92.25
+Appendix J Visualization of the Learned Feature Space
 To illustrate the differences between Max Suppression Regularization and Label Smoothing, we follow the projection technique of Müller et al.
-(2019).
+[21].
 Specifically, we select three semantically related classes and construct an orthonormal basis for the plane intersecting their class templates in feature space.
@@
 Observations
-As shown in Figures 5 and 4, models trained with Max Suppression exhibit:
+As shown in Figures 4 and 3, models trained with Max Suppression exhibit:
   • Enhanced inter-class separability. Distinct classes occupy more clearly separated regions, aligning with improved classification performance.
@@
 Logit Penalty 0.2840 0.3144 0.7996 0.7909 0.6448 0.6024
+Appendix K Evaluation of Calibration Performance
+We have additionally evaluated the effect of MaxSup on calibration in Table 13, based on the Expected Calibration Error (ECE) [8].
+Table 13: Expected Calibration Error (ECE) on CIFAR100. Lower is better.
+ce $10.3\%$
+ls $3.8\%$
+ols $2.9\%$
+zipf $3.5\%$
+maxsup $2.6\%$
+Appendix L Ablation on the Weight Schedule
+We conducted an ablation study on the $\alpha$ schedule, as shown in Table 14.
+The consistently high accuracy across settings demonstrates the robustness of MaxSup.
+The adaptive $\alpha$ schedule, adopted from [18], further highlights the method’s integrity and compatibility with principled design choices.
+Table 14: Ablation study on alpha schedules using ResNet50 on ImageNet1K.
+Schedule $\alpha=0.1+0.1\,\frac{t}{T}$ $\alpha=0+0.1\,\frac{t}{T}$ $\alpha=0.2+0.1\,\frac{t}{T}$ Adaptive Alpha
+77.65 77.62 77.43 77.70
+Appendix M Analysis of Computation Efficiency
+Beyond the standard cross-entropy operations, MaxSup only requires:
+  1. 1.
+    A max operation to determine the largest logit ($O(K)$ complexity),
+  2. 2.
+    A mean operation over the $K$-dimensional logit vector, and
+  3. 3.
+    One subtraction between these two scalars.
+Since $K$ (the number of classes) is usually small (e.g., 1000 for ImageNet-1K), this overhead is minimal compared to the overall forward/backward pass of a deep network.
+In Table 15, we report the average training time per epoch using a ResNet-50 model on the ImageNet-1K dataset.
+Table 15: Average training time per epoch on ImageNet-1K with ResNet-50.
+Method CE (One-Hot) CE + LS CE + MaxSup
+Time/Epoch 3 min 51 s 3 min 52 s 3 min 51 s
+As seen above, the measured run times are nearly identical across all three configurations.
+Thus, the additional cost of MaxSup is negligible compared to the total computation for large-scale training.
+Appendix N Logits Visualization
+As mentioned in Section 4.1.2, all Label Smoothing variants apply a different penalty on the logits.
+To illustrate the impact of different methods on the logits, we plot the histogram of logits of ResNet-50 networks trained with each method over the ImageNet validation set, as shown in Figure 5.
+Figure 5: Comparison of logit distributions under different regularizers.

### v3

@@
 Through extensive feature-space analyses, we show that MaxSup restores intra-class variation and sharpens inter-class boundaries.
-Experiments on large-scale image classification and multiple downstream tasks confirm that MaxSup is a more robust alternative to LS, consistently reducing overconfidence while preserving richer feature representations.
-[^1] Code is available at: https://anonymous.4open.science/r/Maximum-Suppression-Regularization-DB0C.
+Experiments on large-scale image classification and multiple downstream tasks confirm that MaxSup is a more robust alternative to LS.
+[^1] https://github.com/ZhouYuxuanYX/Maximum-Suppression-Regularization.
 1 Introduction
-Multi-class classification [17, 24] typically relies on one-hot labels, which implicitly treat different classes as mutually orthogonal.
-In practice, however, classes often share low-level features [29, 40] or exhibit high-level semantic similarities [3, 22, 38], rendering the one-hot assumption overly restrictive.
-Such a mismatch can yield over-confident classifiers and ultimately degrade generalization [10].
-To moderate overconfidence, Szegedy et al.
-[32] introduced Label Smoothing (LS), which combines a uniform distribution with the hard ground-truth label, thereby reducing the model’s certainty in the primary class.
-LS has since become prevalent in image recognition [11, 20, 33, 43] and neural machine translation [1, 6], often boosting accuracy and calibration [21].
-Yet subsequent work indicates that LS can overly compress features into tight clusters [13, 26, 37], hindering intra-class variability and transferability [5].
-In parallel, Zhu et al.
-[44] found that LS paradoxically fosters overconfidence in misclassified samples, though the precise mechanism behind this remains uncertain.
+Multi-class classification [lecun1998mnist, russakovsky2015imagenet] typically relies on one-hot labels, which implicitly treat different classes as mutually orthogonal.
+In practice, however, classes often share low-level features [silla2011survey, zeiler2014visualizing] or exhibit high-level semantic similarities [chen2021hsva, novack2023chils, yi2022exploring], rendering the one-hot assumption overly restrictive.
+Such a mismatch can yield over-confident classifiers and ultimately degrade generalization [guo2020online].
+To moderate overconfidence, szegedy2016rethinking introduced Label Smoothing (LS), which combines a uniform distribution with the hard ground-truth label, thereby reducing the model’s certainty in the primary class.
+LS has since become prevalent in image recognition [he2016deep, liu2021swin, touvron2021training, zhou2022sp] and neural machine translation [alves2023steering, gao2020towards], often boosting accuracy and calibration [muller2019does].
+Yet subsequent work indicates that LS can overly compress features into tight clusters [kornblith2021better, sariyildiz2022no, xu2023quantifying], hindering intra-class variability and transferability [feng2021rethinking].
+In parallel, zhu2022rethinking found that LS paradoxically fosters overconfidence in misclassified samples, though the precise mechanism behind this remains uncertain.
 In this paper, we reveal that LS’s training objective inherently contains an error amplification term.
-This term pushes the network to reinforce incorrect predictions with exaggerated certainty, yielding highly confident misclassifications and further compressing feature clusters (Section 3.1, Table 1).
-Building on Zhu et al.
-[44], we characterize “overconfidence” in terms of the model’s top-1 prediction, rather than through conventional calibration metrics.
-Through our analysis, we further show that punishing the ground-truth logit during misclassification reduces intra-class variation (Table 2), a phenomenon corroborated by Grad-CAM visualizations (Figure 2).
+This term pushes the network to reinforce incorrect predictions with exaggerated certainty, yielding highly confident misclassifications and further compressing feature clusters (Section˜3.1, Table˜1).
+Building on zhu2022rethinking, we characterize “overconfidence” in terms of the model’s top-1 prediction, rather than through conventional calibration metrics.
+Through our analysis, we further show that punishing the ground-truth logit during misclassification reduces intra-class variation (Table˜2), a phenomenon corroborated by Grad-CAM visualizations (Figure˜2).
 To overcome these shortcomings, we introduce Max Suppression (MaxSup), a method that retains the beneficial regularization effect of LS while eliminating its error amplification.
@@
 By preserving the ground-truth logit in misclassifications, MaxSup sustains richer intra-class variability and sharpens inter-class boundaries.
-As visualized in Figure 1, this approach mitigates the feature collapse and attention drift often induced by LS, ultimately leading to more robust representations.
-Through comprehensive experiments in both image classification (Section 4.2) and semantic segmentation (Section 4.3), we show that MaxSup not only alleviates severe intra-class collapse but also consistently boosts top-1 accuracy (+0.4%–1.0% improvement over LS on ImageNet) and robustly enhances downstream transfer performance (Section 4.1).
+As visualized in Figure˜1, this approach mitigates the feature collapse and attention drift often induced by LS, ultimately leading to more robust representations.
+Through comprehensive experiments in both image classification (Section 4.2) and semantic segmentation (Section 4.3), we show that MaxSup not only alleviates severe intra-class collapse but also consistently boosts top-1 accuracy and robustly enhances downstream transfer performance (Section 4.1).
 Our contributions are summarized as follows:
@@
 As shown in extensive ablations, MaxSup alleviates intra-class collapse and yields consistent accuracy gains.
-  • We demonstrate superior performance across tasks and architectures, including ResNet, MobileNetV2, and DeiT-S, where MaxSup significantly boosts accuracy on ImageNet (up to +1.0% in top-1) and consistently delivers stronger representations for downstream tasks such as semantic segmentation and robust transfer learning.
+  • We demonstrate superior performance across tasks and architectures, including ResNet, MobileNetV2, and DeiT-S, where MaxSup significantly boosts accuracy on ImageNet and consistently delivers stronger representations for downstream tasks such as semantic segmentation and robust transfer learning.
 Figure 1: Comparison of Label Smoothing (LS) and MaxSup.
@@
 Regularization techniques aim to improve the generalization of deep neural networks by constraining model complexity.
-Classical methods like $\ell_{2}$ [16] and $\ell_{1}$ [45] impose direct penalties on large or sparse weights, while Dropout [30] randomly deactivates neurons to discourage over-adaptation.
-In the realm of loss-based strategies, Label Smoothing (LS) [32] redistributes a fraction of the label probability mass away from the ground-truth class, thereby improving accuracy and calibration [21].
-Variants such as Online Label Smoothing (OLS) [41] and Zipf Label Smoothing (Zipf-LS) [19] refine LS by dynamically adjusting the smoothed labels based on a model’s evolving predictions.
-However, they do not fully address the fundamental issue that emerges when the ground-truth logit is not the highest one (see Section 3.1, Table 1).
+Classical methods like $\ell_{2}$ [krogh1991simple] and $\ell_{1}$ [zou2005regularization] impose direct penalties on large or sparse weights, while Dropout [Srivastava2014] randomly deactivates neurons to discourage over-adaptation.
+In the realm of loss-based strategies, Label Smoothing (LS) [szegedy2016rethinking] redistributes a fraction of the label probability mass away from the ground-truth class, thereby improving accuracy and calibration [muller2019does].
+Variants such as Online Label Smoothing (OLS) [zhang2021delving] and Zipf Label Smoothing (Zipf-LS) [liang2022efficient] refine LS by dynamically adjusting the smoothed labels based on a model’s evolving predictions.
+However, they do not fully address the fundamental issue that emerges when the ground-truth logit is not the highest one (see Section˜3.1, Table 1).
 Other loss-based regularizers focus on alternative aspects of the predictive distribution.
-Confidence Penalty [23] penalizes the model’s confidence directly, while Logit Penalty [4] minimizes the global $\ell_{2}$-norm of logits, a technique reported to enhance class separation [13].
-Despite these benefits, Logit Penalty can inadvertently shrink intra-class variation, thereby hampering transfer learning (see Section 4.1).
+Confidence Penalty [pereyra2017regularizing] penalizes the model’s confidence directly, while Logit Penalty [dauphin2021deconstructing] minimizes the global $\ell_{2}$-norm of logits, a technique reported to enhance class separation [kornblith2021better].
+Despite these benefits, Logit Penalty can inadvertently shrink intra-class variation, thereby hampering transfer learning (see Section˜4.1).
 Unlike the aforementioned methods, MaxSup enforces regularization by penalizing only the top-1 logit $z_{\textit{max}}$ rather than the ground-truth logit $z_{gt}$.
@@
 Label Smoothing has also been studied extensively under knowledge distillation.
-For instance, Yuan et al.
-[39] observed that LS can approximate the effect of a teacher–student framework, while Shen et al.
-[28] investigated its role in such pipelines more systematically.
-Additionally, Chandrasegaran et al.
-[2] demonstrated that a low-temperature, LS-trained teacher can notably improve distillation outcomes.
-Concurrently, Kornblith et al.
-[13] showed that LS tightens intra-class clusters in the feature space, diminishing transfer performance.
-From a Neural Collapse perspective [42, 9], LS nudges the model toward rigid feature clusters, as evidenced by the reduced feature variability measured in Xu and Liu [37].
+For instance, yuan2020revisiting observed that LS can approximate the effect of a teacher–student framework, while shen2021label investigated its role in such pipelines more systematically.
+Additionally, chandrasegaran2022revisiting demonstrated that a low-temperature, LS-trained teacher can notably improve distillation outcomes.
+Concurrently, kornblith2021better showed that LS tightens intra-class clusters in the feature space, diminishing transfer performance.
+From a Neural Collapse perspective [zhou2022all, guo2024cross], LS nudges the model toward rigid feature clusters, as evidenced by the reduced feature variability measured in xu2023quantifying.
 Our goal is to overcome LS’s inherent error amplification effect.
 Rather than adjusting how the smoothed label distribution is constructed (as in OLS or Zipf-LS), MaxSup directly penalizes the highest logit $z_{\textit{max}}$.
-This design ensures consistent regularization even if $z_{gt}$ is not the top logit, thereby avoiding the degradation in performance typical of misclassified samples under LS (see Section 3.2).
+This design ensures consistent regularization even if $z_{gt}$ is not the top logit, thereby avoiding the degradation in performance typical of misclassified samples under LS (see Section˜3.2).
 Moreover, MaxSup integrates seamlessly into standard training pipelines, introducing negligible computational overhead beyond substituting the LS term.
@@
 We first partition the training objective into two components: the standard Cross-Entropy (CE) loss and a regularization term introduced by Label Smoothing (LS).
-By expressing LS in terms of logits (Theorem 3.3), we isolate two key factors: a regularization term that controls overconfidence and an error amplification term that enlarges the gap between the ground-truth logit $z_{gt}$ and any higher logits (Corollary 3.4, Equation 5), ultimately degrading performance.
-To address these issues, we propose Max Suppression Regularization (MaxSup), which applies the penalty to the largest logit $z_{\textit{max}}$ rather than $z_{gt}$ (Equation 8, Section 3.2).
+By expressing LS in terms of logits (Theorem˜3.3), we isolate two key factors: a regularization term that controls overconfidence and an error amplification term that enlarges the gap between the ground-truth logit $z_{gt}$ and any higher logits (Corollary˜3.4, Equation˜5), ultimately degrading performance.
+To address these issues, we propose Max Suppression Regularization (MaxSup), which applies the penalty to the largest logit $z_{\textit{max}}$ rather than $z_{gt}$ (Equation˜8, Section˜3.2).
 This shift delivers consistent regularization for both correct and incorrect predictions, preserves intra-class variation, and bolsters inter-class separability.
-Consequently, MaxSup mitigates the representation collapse found in LS, attains superior ImageNet-1K accuracy (Table 1), and improves transferability (Table 2, Table 3).
+Consequently, MaxSup mitigates the representation collapse found in LS, attains superior ImageNet-1K accuracy (Table˜1), and improves transferability (Table˜2, Table˜3).
 The following sections elaborate on MaxSup’s formulation and integration into the training pipeline.
@@
 Label Smoothing (LS) is a regularization technique designed to reduce overconfidence by softening the target distribution.
-Rather than assigning probability $1$ to the ground-truth class and 00 to all others, LS redistributes a fraction $\alpha$ of the probability uniformly across all classes:
+Rather than assigning probability $1$ to the ground-truth class and 0 to all others, LS redistributes a fraction $\alpha$ of the probability uniformly across all classes:
 Definition 3.1.
@@
 This shows that LS adds a regularization term, $L_{\textit{LS}}$, which smooths the output distribution and helps to reduce overfitting.
-(See Appendix A for a formal proof.)
-Building on Lemma 3.2, we next explicitly express $L_{\textit{LS}}$ at the logit level for further analysis.
+(See Appendix˜B for a formal proof.)
+Building on Lemma˜3.2, we next explicitly express $L_{\textit{LS}}$ at the logit level for further analysis.
 Theorem 3.3.
@@
 Thus, LS penalizes the gap between $z_{gt}$ and the average logit, encouraging a more balanced output distribution and reducing overconfidence.
-(See Appendix B for the proof.)
+(See Appendix˜C for the proof.)
 The behavior of $L_{\textit{LS}}$ differs depending on whether $z_{gt}$ is already the maximum logit.
@@
 $L_{\textit{LS}}\;=\;\underbrace{\frac{\alpha}{K}\sum_{z_{m}<z_{gt}}\Bigl(z_{gt}-z_{m}\Bigr)}_{\text{Regularization}}\;+\;\underbrace{\frac{\alpha}{K}\sum_{z_{n}>z_{gt}}\Bigl(z_{gt}-z_{n}\Bigr)}_{\text{Error amplification}},$ (5)
-where $M$ and $N$ are the numbers of logits below and above $z_{gt}$, respectively ($M+N=K-1$). Note that the error amplification term vanishes when $z_{gt}=z_{\textit{max}}$.
+where $M$ and $N$ are the numbers of logits below and above $z_{gt}$, respectively (M+N=K1M+N=K-1). Note that the error amplification term vanishes when $z_{gt}=z_{\textit{max}}$.
   1. 1.
@@
     • Inconsistent Regularization: The regularization term lowers $z_{gt}$ yet does not penalize $z_{\textit{max}}$, which further impairs learning.
-These issues with LS on misclassified samples have also been systematically observed in prior work [35].
+These issues with LS on misclassified samples have also been systematically observed in prior work [xia2024understanding].
 By precisely disentangling these two components (regularization vs.
@@
 + Label Smoothing $\tfrac{\alpha}{K}\sum_{z_{m}<z_{gt}}(z_{gt}-z_{m})$ 75.91
-$\,+\,\tfrac{\alpha}{K}\sum_{z_{n}>z_{gt}}(z_{gt}-z_{n})$
+$\tfrac{\alpha}{K}\sum_{z_{n}>z_{gt}}(z_{gt}-z_{n})$
 + Regularization $\tfrac{\alpha}{M}\sum_{z_{m}<z_{gt}}(z_{gt}-z_{m})$ 75.98
@@
 Ablation Study on LS Components.
-To isolate the effects of each component in LS, we carefully perform a detailed and systematic ablation study on ImageNet-1K using a DeiT-Small model [33] without Mixup or CutMix.
-As indicated in Table 1, the performance gains from LS stem solely from the regularization term, whereas the error amplification term degrades accuracy.
+To isolate the effects of each component in LS, we carefully perform a detailed and systematic ablation study on ImageNet-1K using a DeiT-Small model [touvron2021training] without Mixup or CutMix.
+As indicated in Table˜1, the performance gains from LS stem solely from the regularization term, whereas the error amplification term degrades accuracy.
 In contrast, our MaxSup omits the error amplification component and leverages only the beneficial regularization, thereby boosting accuracy beyond that of standard LS.
-Specifically, Table 1 shows that LS’s overall improvement can be attributed exclusively to its regularization contribution; the error amplification term consistently reduces accuracy (e.g., to 73.63% or 73.69%).
+Specifically, Table˜1 shows that LS’s overall improvement can be attributed exclusively to its regularization contribution; the error amplification term consistently reduces accuracy (e.g., to 73.63% or 73.69%).
 Disabling only the error amplification while retaining the regularization yields a slight but measurable improvement (75.98% vs.
@@
 3.2 Max Suppression Regularization
-Building on our analysis in Section 3.1, we find that Label Smoothing (LS) not only impacts correctly classified samples but also influences misclassifications in unintended and harmful ways.
+Building on our analysis in Section˜3.1, we find that Label Smoothing (LS) not only impacts correctly classified samples but also influences misclassifications in unintended and harmful ways.
 Specifically, LS suffers from two main limitations: inconsistent regularization and error amplification.
@@
 We define the Cross-Entropy loss with MaxSup as follows:
-$\underbrace{H(\mathbf{s},\mathbf{q})}_{\text{CE with Soft Labels}}\;=\;\underbrace{H(\mathbf{y},\mathbf{q})}_{\text{CE with Hard Labels}}\;+\;\underbrace{L_{\textit{MaxSup}}}_{\text{Max Suppression Loss}},$ (6)
+[…]
 where
@@
 and
-[…]
+$y^{\prime}_{k}\;=\;\mathds{1}_{\bigl\{\,k=\arg\max(\mathbf{q})\bigr\}},$
 so that $y^{\prime}_{k}=1$ identifies the model’s top-1 prediction and $y^{\prime}_{k}=0$ otherwise.
@@
 By shifting the penalty from $z_{gt}$ (the ground-truth logit) to $z_{\textit{max}}$ (the highest logit), MaxSup avoids unduly suppressing $z_{gt}$ when the model misclassifies, thus overcoming Label Smoothing’s principal shortcoming.
-Logit-Level Formulation of MaxSup. Building on the logit-level perspective introduced for LS in Section 3.1, we can express $L_{\textit{MaxSup}}$ as:
+Logit-Level Formulation of MaxSup. Building on the logit-level perspective introduced for LS in Section˜3.1, we can express $L_{\textit{MaxSup}}$ as:
 $L_{\textit{MaxSup}}\;=\;\alpha\Bigl(z_{\textit{max}}\;-\;\tfrac{1}{K}\sum_{k=1}^{K}z_{k}\Bigr),$ (8)
@@
 However, when $z_{gt}\neq z_{\textit{max}}$, LS shrinks $z_{gt}$, widening the gap with the incorrect logit, whereas MaxSup penalizes $z_{\textit{max}}$, preserving $z_{gt}$ from undue suppression.
-As illustrated in Figure 2, this helps the model recover from mistakes more effectively and avoid reinforcing incorrect predictions.
+As illustrated in Figure˜2, this helps the model recover from mistakes more effectively and avoid reinforcing incorrect predictions.
 Gradient Analysis. To understand MaxSup’s optimization dynamics, we compute its gradients with respect to each logit $z_{k}$. Specifically,
@@
 In misclassified cases, the ground-truth logit $z_{gt}$ is spared from penalization, avoiding the error-amplification issue seen in LS.
-For completeness, Appendix A provides the full gradient derivation.
+For completeness, Appendix B provides the full gradient derivation.
+While [xia2024understanding] conducted a related gradient analysis of the training loss, it focuses specifically on the setting of selective classification, and examines a posthoc logit normalization technique to mitigate confidence calibration issues.
+However, this approach addresses only the overconfidence problem of label smoothing (LS), without tackling representation collapse.
+Moreover, our work presents a logit-level reformulation of LS that provides a deeper theoretical understanding of why LS amplifies errors.
 Behavior Across Different Samples.
@@
 4.1 Analysis of MaxSup’s Learning Benefits
-Having established how MaxSup addresses Label Smoothing’s (LS) principal shortcomings (Section 3.1), we now demonstrate its impact on inter-class separability and intra-class variation—two properties essential for accurate classification and effective transfer learning.
-Table 2: Feature quality metrics for ResNet-50 on ImageNet-1K.
-While all methods reduce $\bar{d}_{\text{within}}$ (intra-class variation) relative to the baseline, MaxSup retains the greatest within-class diversity.
-Both $\bar{d}_{\text{within}}$ and $R^{2}$ benefit from higher values.
+Having established how MaxSup addresses Label Smoothing’s (LS) principal shortcomings (Section˜3.1), we now demonstrate its impact on inter-class separability and intra-class variation—two properties essential for accurate classification and effective transfer learning.
+Table 2: Feature quality of ResNet-50 on ImageNet-1K.
 Method $\bar{d}_{\text{within}}\!\uparrow$ $R^{2}\!\uparrow$
@@
 By contrast, MaxSup uniformly penalizes the top-1 logit, whether the prediction is correct or incorrect, thereby eliminating LS’s error-amplification effect and preserving finer distinctions within each class.
-Table 2 compares intra-class variation ($\bar{d}_{\text{within}}$) and inter-class separability ($R^{2}$[13] for ResNet-50 trained on ImageNet-1K.
+Table 2 compares intra-class variation ($\bar{d}_{\text{within}}$) and inter-class separability ($R^{2}$[kornblith2021better] for ResNet-50 trained on ImageNet-1K.
 Although all investigated regularizers decrease $\bar{d}_{\text{within}}$ relative to a baseline, MaxSup yields the smallest reduction, indicating a stronger retention of subtle within-class diversity—widely associated with enhanced generalization and improved transfer performance.
@@
 By contrast, MaxSup preserves near-baseline performance, implying that it maintains rich discriminative information crucial for downstream tasks.
-For extended evaluations on diverse datasets, see Table 9 in the appendix.
+For extended evaluations on diverse datasets, see Table˜12 in the appendix.
 Table 3: Linear-probe transfer accuracy on CIFAR-10 (higher is better).
@@
 4.1.2 Connection to Logit Penalty
-As detailed in Section 3, both Label Smoothing (LS) variants and MaxSup impose penalties directly at the logit level, aligning with the perspective that various regularizers influence a model’s representational capacity via distinct logit constraints [13].
+As detailed in Section 3, both Label Smoothing (LS) variants and MaxSup impose penalties directly at the logit level, aligning with the perspective that various regularizers influence a model’s representational capacity via distinct logit constraints [kornblith2021better].
 Within this family of techniques, Logit Penalty and MaxSup both address the maximum logit, yet diverge fundamentally in their specific methods of regularization.
@@
 Because only the highest-confidence prediction is penalized, MaxSup avoids the uniform shrinkage observed in Logit Penalty, preserving richer intra-class variation—a property essential for robust transfer.
-Further insights into this behavior can be found in Appendix N, where logit-value histograms illustrate how each method affects the logit distribution.
+Further insights into this behavior can be found in Appendix˜M, where logit-value histograms illustrate how each method affects the logit distribution.
 4.2 Evaluation on ImageNet Classification
@@
 4.2.1 Experiment Setup
-Model Training Configurations.We evaluate both convolutional (ResNet[11], MobileNetV2 [25]) and transformer (DeiT-Small [33]) architectures on ImageNet [15].
+Model Training Configurations.We evaluate both convolutional (ResNet[he2016deep], MobileNetV2 [sandler2018mobilenetv2]) and transformer (DeiT-Small [touvron2021training]) architectures on ImageNet [krizhevsky2012imagenet].
 For the ResNet Series, we train for 200 epochs using stochastic gradient descent (SGD) with momentum0.9, weight decay of $1\times 10^{-4}$, and a batch size of 2048.
@@
 [^2] Additional training hyperparameters follow the FFCV scripts at https://github.com/libffcv/ffcv.
-See Appendix E for further details.
+See Appendix F for further details.
 We also test ResNet variants on CIFAR-100 with a conventional setup: an initial learning rate of 0.1 (reduced fivefold at epochs 60, 120, and 160), training for 200 epochs with batch size 128 and weight decay $5\times 10^{-4}$.
-For DeiT-Small, we use the official codebase [33], training from scratch without knowledge distillation to isolate MaxSup’s contribution.
+For DeiT-Small, we use the official codebase [touvron2021training], training from scratch without knowledge distillation to isolate MaxSup’s contribution.
 CutMix and Mixup are disabled to ensure the model optimization objective remains unchanged.
-Hyperparameters for Compared Methods.We compare Max Suppression Regularization against a range of LS extensions, including Zipf Label Smoothing[19] and Online Label Smoothing [41].
+Hyperparameters for Compared Methods.We compare Max Suppression Regularization against a range of LS extensions, including Zipf Label Smoothing[liang2022efficient] and Online Label Smoothing [zhang2021delving].
 Where official implementations exist, we adopt them directly; otherwise, we follow the methodological details provided in each respective paper.
 Except for any method-specific hyperparameters, all other core training settings remain identical to the baselines.
-Furthermore, both MaxSup and standard LS employ a linearly increasing $\alpha$-scheduler for improved training stability (see Appendix F).
+Furthermore, both MaxSup and standard LS employ a linearly increasing $\alpha$-scheduler for improved training stability (see Appendix˜G).
 This ensures a fair comparison under consistent and reproducible training protocols.
@@
 Across all convolutional architectures tested, MaxSup consistently delivers the highest top-1 accuracy among label-smoothing approaches.
-By contrast, OLS [41] and Zipf-LS [19] exhibit less stable gains, suggesting their effectiveness may heavily hinge on specific training protocols.
+By contrast, OLS [zhang2021delving] and Zipf-LS [liang2022efficient] exhibit less stable gains, suggesting their effectiveness may heavily hinge on specific training protocols.
 To reproduce OLS and Zipf-LS, we apply the authors’ official codebases and hyperparameters but do not replicate their complete training recipes (e.g., OLS trains for 250 epochs with a step-scheduled learning rate of 0.1, and Zipf-LS uses 100 epochs with distinct hyperparameters).
@@
 Label Smoothing 69.54±0.15 76.91±0.11 77.37±0.15 71.61±0.09 77.05±0.17 78.88±0.13 79.19±0.25 69.65±0.08
-[…]
+$\text{Zipf-LS}^{*}$ 69.31±0.12 76.73±0.17 76.91±0.11 71.16±0.15 76.21±0.12 78.75±0.21 79.15±0.18 69.39±0.08
 $\text{OLS}^{*}$ 69.45±0.15 77.23±0.21 77.71±0.17 71.63±0.11 77.33±0.15 78.79±0.12 79.25±0.15 68.91±0.11
@@
 Baseline 74.39 0.19
-Label Smoothing 76.08 ($+1.69$) 0.16
-Zipf-LS 75.89 ($+1.50$) 0.26
-OLS 76.16 ($+1.77$) 0.18
-MaxSup 76.49 ($+2.10$) 0.12
+Label Smoothing 76.08 (+1.69+1.69) 0.16
+Zipf-LS 75.89 (+1.50+1.50) 0.26
+OLS 76.16 (+1.77+1.77) 0.18
+MaxSup 76.49 (+2.10+2.10) 0.12
 DeiT Comparison.
@@
 These findings underscore MaxSup’s adaptability to different architectures and emphasize its utility in scenarios where conventional label-smoothing methods yield limited benefits.
+Table 6: Classification on CUB and Cars Datasets.
+Method CUB[cub] Cars[car]
+Baseline 80.88 90.27
+LS 81.96 91.64
+OLS 82.33 91.96
+Zipf-LS 81.40 90.99
+MaxSup 82.53 92.25
 Fine-Grained Classification.
-Beyond ImageNet, we evaluate MaxSup on demanding fine-grained benchmarks such as CUB-200 [34] and Stanford Cars [14].
-As shown in Table 11, MaxSup outperforms other LS-based methods by more effectively capturing subtle within-class variations, which are critical for success in fine-grained recognition scenarios.
+Beyond large-scale benchmarks like ImageNet, we further evaluate MaxSup on two fine-grained visual recognition tasks: CUB-200-2011 [cub] and Stanford Cars [car].
+These datasets pose unique challenges due to subtle inter-class differences, which often expose the limitations of standard regularization approaches.
+As shown in Table˜6, MaxSup achieves the best performance across both datasets, surpassing LS and its recent variants.
+This demonstrates that MaxSup encourages the model to learn more discriminative and semantically rich representations that better capture fine-grained attributes, such as textures and part-level details.
+The consistent improvements on these benchmarks further validate MaxSup’s capacity to generalize across different visual domains and its potential to enhance robustness in recognition scenarios where nuanced feature understanding is critical.
+Table 7: Comparison of classification performance (%) across imbalance levels for different loss strategies (Focal Loss vs Label Smoothing (LS) vs MaxSup) on the long-tailed CIFAR-10 dataset using Resnet-32.
+Best performances are in bold.
+Dataset Split Imbalance Ratio Method Overall Many Medium Low
+LT CIFAR-10 val 50 Focal Loss 77.4 76.0 89.7 0.0
+Label Smoothing 81.2 81.6 77.0 0.0
+MaxSup 82.1 82.5 78.1 0.0
+LT CIFAR-10 test 50 Focal Loss 76.8 75.3 90.4 0.0
+Label Smoothing 80.5 81.1 75.4 0.0
+MaxSup 81.4 82.3 73.4 0.0
+LT CIFAR-10 val 100 Focal Loss 75.1 71.8 88.3 0.0
+Label Smoothing 76.6 80.6 60.7 0.0
+MaxSup 77.1 80.1 65.1 0.0
+LT CIFAR-10 test 100 Focal Loss 74.7 71.6 87.2 0.0
+Label Smoothing 76.4 80.8 59.0 0.0
+MaxSup 76.4 79.9 62.4 0.0
+Long-Tailed Classification.
+To assess the effectiveness of MaxSup under data imbalance, we performed experiments on the CIFAR-10-LT dataset with imbalance ratios of 50 and 100, following the experimental settings described in [tang2020long].
+The corresponding results are summarized in Table˜7.
+The evaluation compares three setups: Focal Loss, Focal Loss + LS, and Focal Loss + MaxSup.
+Across all imbalance ratios and splits (val/test), MaxSup consistently outperforms both the baseline and LS in overall accuracy, which jointly reflects the many-shot, medium-shot, and low-shot (minor class) performance.
+For example, at an imbalance ratio of 50 on the test split, MaxSup achieves $81.4\%$ accuracy, outperforming Focal Loss ($76.8\%$) by 4.6 percentage points, and LS ($80.5\%$) by 0.9 percentage points.
+These results indicate that MaxSup achieves a better trade-off between many- and medium-shot accuracy.
+While it does not fully resolve the challenge of imbalanced classification—especially for minority classes—it shows positive effects and offers a promising direction for further extension.
+Table 8: Comparison of MaxSup, Label Smoothing (LS), and standard Cross Entropy (CE) on CIFAR-10-C. Lower is better. Values show mean(std) across three setups.
+Metric MaxSup LS CE
+Error (Corr) 0.362(0.055) 0.359(0.064) 0.354(0.015)
+NLL (Corr) 1.770(0.103) 1.476(0.111) 1.819(0.158)
+ECE (Corr) 0.145(0.003) 0.158(0.015) 0.260(0.015)
+Corrupted Image Classification To evaluate the effectiveness of MaxSup on out-of-distribution (OOD) settings, we also conducted experiments on CIFAR10-C benchmark [hendrycksbenchmarking] shown in Table˜8 following settings in [heinonen2025robust].
+Table˜8 reports the performance of MaxSup and Label Smoothing (LS) on this benchmark using ResNet-50 as the backbone.
+Specifically, LS yields a better NLL (1.5730 vs.
+1.8431), implying more confident probabilistic predictions.
+However, MaxSup achieves a better ECE (0.1479 vs.
+0.1741), indicating better calibration of the predicted confidence scores.
+These results validate that MaxSup remains effective on OOD datasets, achieving performance comparable to LS across all three metrics.
 Ablation on the Weight Schedule.
@@
 Empirical results indicate that MaxSup consistently maintains high accuracy across a wide range of schedules, further underscoring its robustness against hyperparameter changes.
-For additional details and discussions, refer to Appendix F.
+For additional details and discussions, refer to Appendix˜G.
 Table 9: Semantic segmentation (multi-scale) on ADE20K using UperNet. All models are pretrained on ImageNet-1K; mIoU reported as percentage.
@@
 DeiT-Small Baseline 42.1
-Label Smoothing 42.4 ($+0.3$)
-MaxSup 42.8 ($+0.7$)
+Label Smoothing 42.4 (+0.3+0.3)
+MaxSup 42.8 (+0.7+0.7)
 4.3 Evaluation on Semantic Segmentation
 We further investigate MaxSup’s applicability to downstream tasks by evaluating its performance on semantic segmentation using the widely adopted MMSegmentation framework.
-[^3] https://github.com/open-mmlab/mmsegmentation Specifically, we adopt the UperNet [36] architecture with a DeiT-Small backbone, trained on ADE20K.
-Models pretrained on ImageNet-1K with either MaxSup or Label Smoothing are then fine-tuned under the same cross-entropy objective (Section 4.2.2).
+[^3] https://github.com/open-mmlab/mmsegmentation Specifically, we adopt the UperNet [xiao2018unified] architecture with a DeiT-Small backbone, trained on ADE20K.
+Models pretrained on ImageNet-1K with either MaxSup or Label Smoothing are then fine-tuned under the same cross-entropy objective (Section˜4.2.2).
 Table 9 shows that initializing with MaxSup-pretrained weights yields an mIoU of $42.8\%$, surpassing the $42.4\%$ achieved by Label Smoothing.
@@
 By more effectively capturing class boundaries and within-class variability, MaxSup promotes stronger segmentation results, underscoring its potential to deliver features that are both transferable and highly robust.
-Figure 2: Grad-CAM [27] visualizations for DeiT-Small models under three training setups: MaxSup (2nd row), Label Smoothing (3rd row), and a baseline (4th row).
+Figure 2: Grad-CAM [Selvaraju_2019] visualizations for DeiT-Small models under three training setups: MaxSup (2nd row), Label Smoothing (3rd row), and a baseline (4th row).
 The first row shows the original images.
@@
 4.4 Visualization via Class Activation Maps
-To better understand how MaxSup fundamentally differs from Label Smoothing (LS) in guiding model decisions, we employ Gradient-weighted Class Activation Mapping (Grad-CAM) [27], which highlights regions most influential for each prediction.
+To better understand how MaxSup fundamentally differs from Label Smoothing (LS) in guiding model decisions, we employ Gradient-weighted Class Activation Mapping (Grad-CAM) [Selvaraju_2019], which highlights regions most influential for each prediction.
 We evaluate DeiT-Small under three training setups: MaxSup (second row), LS (third row), and a baseline with standard cross-entropy (fourth row).
-As illustrated in Figure 2, MaxSup-trained models more effectively suppress background distractions than LS, which often fixates on unrelated objects—such as poles in “Bird,” tubes in “Goldfish,” and caps in “House Finch.” This behavior reflects LS’s error-enhancement mechanism, which can misdirect attention.
+As illustrated in Figure˜2, MaxSup-trained models more effectively suppress background distractions than LS, which often fixates on unrelated objects—such as poles in “Bird,” tubes in “Goldfish,” and caps in “House Finch.” This behavior reflects LS’s error-enhancement mechanism, which can misdirect attention.
 Moreover, MaxSup retains a wider spectrum of salient features, as exemplified in the Shark” and Monkey” images, where LS-trained models often omit crucial semantic details (e.g., fins, tails, or facial contours).
-These findings align with our analysis in Appendix J, clearly demonstrating that MaxSup preserves richer intra-class information.
+These findings align with our analysis in Appendix˜J, clearly demonstrating that MaxSup preserves richer intra-class information.
 Consequently, MaxSup-trained models produce more accurate and consistent predictions by effectively leveraging fine-grained object cues.
@@
 Limitations.
-Although promising, several questions remain.
-Prior work[21] notes that LS-trained teachers may degrade knowledge distillation [12], and Guo et al.
-[9] suggests LS accelerates convergence via conditioning.
-Examining MaxSup’s role in distillation and its impact on training dynamics would clarify these effects.
-Recent studies [31, 7] also show that $\ell_{2}$ regularization biases final-layer features toward low-rank solutions, raising questions about whether MaxSup behaves similarly.
+Prior work [muller2019does] notes that LS-trained teachers may degrade knowledge distillation [hinton2015distilling, Hu2021CVPR], and guo2024cross suggests LS accelerates convergence via improved conditioning.
+Examining MaxSup’s potential role in distillation and its overall impact on training dynamics would clarify these underlying effects.
+Recent studies [sukenik2024neural, garrod2024persistence] also show that $\ell_{2}$ regularization biases final-layer features toward low-rank solutions, raising interesting questions about whether MaxSup behaves similarly.
 Impact.
@@
 By offering researchers and practitioners both a clearer understanding of LS’s limitations and a straightforward, computationally light, and easily integrable method to overcome them, MaxSup may help guide the development of more reliable and interpretable deep learning models.
+Acknowledgments
+This work was supported by the University of Washington Faculty Startup Fund, the Carwein–Andrews Fellowship, the UW GSFEI Top Scholar Award, and the U.S.
+DOT PacTrans sub-center seed funding program.
+We also thank the anonymous reviewers for their constructive feedback and insightful suggestions, which helped improve the quality and clarity of this paper.
 NeurIPS Paper Checklist
@@
 The LLM was not involved in designing, conducting, or analyzing the experiments, nor in developing the core algorithmic contributions.
-Appendix A Proof of Lemma 3.2
+Appendix A Technical Appendices and Supplementary Material
+Technical appendices with additional results, figures, graphs and proofs may be submitted with the paper submission before the full submission deadline (see above), or as a separate PDF in the ZIP file below before the supplementary material deadline.
+There is no page limit for the technical appendices.
+Appendix B Proof of Lemma 3.2
 Proof.
@@
 It is noteworthy that the original cross-entropy loss $H(\mathbf{y},\mathbf{q})$ remains unweighted by $\alpha$ in this decomposition, which is consistent with the statement in Lemma 3.2 ∎
-Appendix B Proof of Theorem 3.3
+Appendix C Proof of Theorem˜3.3
 Proof.
@@
 This completes the proof for both cases of Theorem 3.3.
-Appendix C Gradient Analysis
-C.1 New Objective Function
+Appendix D Gradient Analysis
+D.1 New Objective Function
 The Cross Entropy with Max Suppression is defined as:
@@
 where $H(\cdot,\cdot)$ denotes the cross-entropy function.
-C.2 Gradient Analysis
+D.2 Gradient Analysis
 The gradient of the loss with respect to the logit $z_{i}$ for each class $i$ is derived as:
@@
 14:end for
-Appendix D Pseudo Code
+Appendix E Pseudo Code
 Algorithm 1 presents pseudo code illustrating gradient descent with Max Suppression (MaxSup).
 The main difference from standard Label Smoothing lies in penalizing the highest logit rather than the ground-truth logit.
-Appendix E Robustness Under Different Training Recipes
+Appendix F Robustness Under Different Training Recipes
 We assess MaxSup’s robustness by testing it under a modified training recipe that reduces total training time and alters the learning rate schedule.
@@
 Logit Penalty 66.97±0.11 74.21±0.16 75.17±0.12 70.39±0.14
-Appendix F Increasing Smoothing Weight Schedule
+Appendix G Increasing Smoothing Weight Schedule
 Building on the intuition that a model’s confidence naturally grows as training progresses, we propose a linearly increasing schedule for the smoothing parameter $\alpha$.
@@
 Experimental Evidence
-As shown in Table 8, both Label Smoothing and MaxSup benefit from this $\alpha$ scheduler.
+As shown in Table˜11, both Label Smoothing and MaxSup benefit from this $\alpha$ scheduler.
 For Label Smoothing, accuracy improves from 75.91% to 76.16%, while MaxSup sees a more pronounced gain, from 76.12% to 76.58%.
-This greater improvement for MaxSup ($+0.46\%$) compared to Label Smoothing ($+0.25\%$) corroborates our claim that MaxSup successfully addresses the inconsistent regularization and error-enhancement issues of Label Smoothing during misclassifications.
+This greater improvement for MaxSup ($0.46\%$) compared to Label Smoothing ($0.25\%$) corroborates our claim that MaxSup successfully addresses the inconsistent regularization and error-enhancement issues of Label Smoothing during misclassifications.
 Table 11: Effect of an $\alpha$ scheduler on model performance.
@@
 MaxSup $\alpha\Bigl(z_{\textit{max}}-\tfrac{1}{K}\sum_{k}z_{k}\Bigr)$ 76.12 76.58
-Appendix G Extended Evaluation of Linear Transferability on Different Datasets
-To further demonstrate the substantial improvement in feature representation compared to other methods, we further compare the linear transfer accuracies of different methods on a broader range of datasets in Table 9
+Appendix H Extended Evaluation of Linear Transferability on Different Datasets
+To further demonstrate the substantial improvement in feature representation compared to other methods, we further compare the linear transfer accuracies of different methods on a broader range of datasets in Table˜12
 Table 12: Validation performance of different methods, evaluated using multinomial logistic regression with l2 regularization.
 Although Label Smoothing and OLS improve ImageNet accuracy, they substantially degrade transfer accuracy compared to MaxSup.
-Following [13], we selected from 45 logarithmically spaced values between $10^{-6}$ and $10^{5}$.
+Following [kornblith2021better], we selected from 45 logarithmically spaced values between $10^{-6}$ and $10^{5}$.
 Notably, the search range is larger than the search range used in Table 3, thus leading to higher overall accuracies on CIFAR10.
@@
 MaxSup 91.00 73.93 67.29 88.84 70.94 92.93
-Appendix H Extended Comparison to More Label Smoothing Alternatives
-We have included Confidence Penalty [23] and Adaptive Label Smooothing with Self-Knowledge [18] for an extended comparison in Table 10.
+Appendix I Extended Comparison to More Label Smoothing Alternatives
+We have included Confidence Penalty [pereyra2017regularizing] and Adaptive Label Smooothing with Self-Knowledge [lee2022adaptive] for an extended comparison in Table˜13.
 We follow their recommended hyperparameter settings: weight coefficient is set to 0.1 for confidence penalty, and the checkpoint with the highest validation accuracy checkpoint is treated as teacher for Adaptive Label Smoothing.
-From Table 10, we see that MaxSup outperforms both Confidence Penalty and Adaptive LS on ImageNet with ResNet-50.
-This aligns with our theoretical analysis that selectively penalizing $z_{\mathrm{max}}$ yields a more consistent regularization than penalizing all logits (Confidence Penalty) or dynamically smoothing the label distribution with self knowledge (Adaptive LS).
-Table 10: Comparison of classic convolutional neural networks on ImageNet and CIFAR-100.
-Results are reported as “mean ± std” (percentage).
+To further address the novelty of our work, we additionally compared our method to recent approaches that also identify and aim to fix the issues of Label Smoothing on misclassified samples [gao2020towards, wei2022mitigatingneuralnetworkoverconfidence].
+Specifically, the method proposed in [wei2022mitigatingneuralnetworkoverconfidence] aims to mitigate overconfidence via logit normalization during training.
+With its default hyperparameter settings, it achieves an accuracy of 74.32% on ImageNet with a ResNet-50, which is significantly lower than the 76.91% achieved by standard Label Smoothing.
+This performance aligns with that of Logit Penalty, which similarly minimizes the global $l_{2}$-norm of logits and can struggle to match baseline LS performance.
+We also note that these norm-based methods[gao2020towards, wei2022mitigatingneuralnetworkoverconfidence, dauphin2021deconstructing] are often highly sensitive to hyperparameter choices, which can limit their practical applicability.
+As shown in 13, MaxSup outperforms all these alternatives on ImageNet with ResNet-50.
+This aligns with our theoretical analysis that selectively penalizing $z_{max}$ yields a more consistent and effective regularization than penalizing all logits (Confidence Penalty), dynamically smoothing the label distribution with self-knowledge (Adaptive LS), or applying global norm-based penalties (Logit Penalty, Logit Normalization).
+Table 13: Comparison of classic convolutional neural networks on ImageNet.
+Results are reported as “mean $\pm$ std” (percentage).
 Bold entries highlight the best performance; underlined entries mark the second best.
@@
 ResNet-50
-Baseline 76.41±0.10
-Label Smoothing 76.91±0.11
-[…]
-$\text{OLS}^{*}$ 77.23±0.21
-MaxSup 77.69±0.07
-Logit Penalty 76.73±0.10
-Confidence Penalty 76.58±0.12
-Adaptive Label Smoothing 77.16±0.15
-Appendix I Classification on Fine-Grained Datasets
-We have further compared MaxSup and other methods on CUB-200 and Stanford Cars for fine-grained classification in Table 11.
-The performance of MaxSup again supports its superiority against existing approaches.
-Table 11: Classification on CUB and Cars Datasets.
-Method CUB[34] Cars[14]
-no_ls 80.88 90.27
-ls 81.96 91.64
-ols 82.33 91.96
-zipfs 81.40 90.99
-max_sup 82.53 92.25
+Baseline $76.41\pm 0.10$
+Label Smoothing $76.91\pm 0.11$
+Zipf-LS $76.73\pm 0.17$
+OLS $77.23\pm 0.21$
+MaxSup $77.69\pm 0.07$
+Logit Penalty $76.73\pm 0.10$
+Logit Normalization [3*] 74.32
+Confidence Penalty $76.58\pm 0.12$
+Adaptive Label Smoothing $77.16\pm 0.15$
 Appendix J Visualization of the Learned Feature Space
-To illustrate the differences between Max Suppression Regularization and Label Smoothing, we follow the projection technique of Müller et al.
-[21].
+To illustrate the differences between Max Suppression Regularization and Label Smoothing, we follow the projection technique of muller2019does.
 Specifically, we select three semantically related classes and construct an orthonormal basis for the plane intersecting their class templates in feature space.
@@
     Confusion.
-We identify a class that the Label Smoothing (LS)–trained model frequently misclassifies and select two additional classes involved in those misclassifications (Figure 3(c), Figure 4(c)).
-Conversely, we also examine a scenario where a class under Max Suppression is confused with others, highlighting key differences (Figure 3(d), Figure 4(d)).
+We identify a class that the Label Smoothing (LS)–trained model frequently misclassifies and select two additional classes involved in those misclassifications (Figure˜3(c), Figure˜4(c)).
+Conversely, we also examine a scenario where a class under Max Suppression is confused with others, highlighting key differences (Figure˜3(d), Figure˜4(d)).
 (a) Semantically Similar Classes
@@
 Observations
-As shown in Figures 4 and 3, models trained with Max Suppression exhibit:
+As shown in Figures˜4 and 3, models trained with Max Suppression exhibit:
   • Enhanced inter-class separability. Distinct classes occupy more clearly separated regions, aligning with improved classification performance.
@@
 Logit Penalty 0.2840 0.3144 0.7996 0.7909 0.6448 0.6024
-Appendix K Evaluation of Calibration Performance
-We have additionally evaluated the effect of MaxSup on calibration in Table 13, based on the Expected Calibration Error (ECE) [8].
-Table 13: Expected Calibration Error (ECE) on CIFAR100. Lower is better.
-ce $10.3\%$
-ls $3.8\%$
-ols $2.9\%$
-zipf $3.5\%$
-maxsup $2.6\%$
-Appendix L Ablation on the Weight Schedule
-We conducted an ablation study on the $\alpha$ schedule, as shown in Table 14.
+Appendix K Ablation on the Weight Schedule
+We conducted an ablation study on the $\alpha$ schedule, as shown in Table˜15.
 The consistently high accuracy across settings demonstrates the robustness of MaxSup.
-The adaptive $\alpha$ schedule, adopted from [18], further highlights the method’s integrity and compatibility with principled design choices.
+The adaptive $\alpha$ schedule, adopted from [lee2022adaptive], further highlights the method’s integrity and compatibility with principled design choices.
 Table 15: Ablation study on alpha schedules using ResNet50 on ImageNet1K.
@@
 77.65 77.62 77.43 77.70
-Appendix M Analysis of Computation Efficiency
+Appendix L Analysis of Computation Efficiency
 Beyond the standard cross-entropy operations, MaxSup only requires:
@@
 Since $K$ (the number of classes) is usually small (e.g., 1000 for ImageNet-1K), this overhead is minimal compared to the overall forward/backward pass of a deep network.
-In Table 15, we report the average training time per epoch using a ResNet-50 model on the ImageNet-1K dataset.
+In Table˜16, we report the average training time per epoch using a ResNet-50 model on the ImageNet-1K dataset.
 Table 16: Average training time per epoch on ImageNet-1K with ResNet-50.
@@
 Thus, the additional cost of MaxSup is negligible compared to the total computation for large-scale training.
-Appendix N Logits Visualization
-As mentioned in Section 4.1.2, all Label Smoothing variants apply a different penalty on the logits.
-To illustrate the impact of different methods on the logits, we plot the histogram of logits of ResNet-50 networks trained with each method over the ImageNet validation set, as shown in Figure 5.
+Appendix M Logits Visualization
+As mentioned in Section˜4.1.2, all Label Smoothing variants apply a different penalty on the logits.
+To illustrate the impact of different methods on the logits, we plot the histogram of logits of ResNet-50 networks trained with each method over the ImageNet validation set, as shown in Figure˜5.
 Figure 5: Comparison of logit distributions under different regularizers.

### v4

@@
 1 Introduction
-Multi-class classification [lecun1998mnist, russakovsky2015imagenet] typically relies on one-hot labels, which implicitly treat different classes as mutually orthogonal.
-In practice, however, classes often share low-level features [silla2011survey, zeiler2014visualizing] or exhibit high-level semantic similarities [chen2021hsva, novack2023chils, yi2022exploring], rendering the one-hot assumption overly restrictive.
-Such a mismatch can yield over-confident classifiers and ultimately degrade generalization [guo2020online].
-To moderate overconfidence, szegedy2016rethinking introduced Label Smoothing (LS), which combines a uniform distribution with the hard ground-truth label, thereby reducing the model’s certainty in the primary class.
-LS has since become prevalent in image recognition [he2016deep, liu2021swin, touvron2021training, zhou2022sp] and neural machine translation [alves2023steering, gao2020towards], often boosting accuracy and calibration [muller2019does].
-Yet subsequent work indicates that LS can overly compress features into tight clusters [kornblith2021better, sariyildiz2022no, xu2023quantifying], hindering intra-class variability and transferability [feng2021rethinking].
-In parallel, zhu2022rethinking found that LS paradoxically fosters overconfidence in misclassified samples, though the precise mechanism behind this remains uncertain.
+Multi-class classification [19, 26] typically relies on one-hot labels, which implicitly treat different classes as mutually orthogonal.
+In practice, however, classes often share low-level features [31, 44] or exhibit high-level semantic similarities [3, 24, 42], rendering the one-hot assumption overly restrictive.
+Such a mismatch can yield over-confident classifiers and ultimately degrade generalization [9].
+To moderate overconfidence, Szegedy et al.
+[34] introduced Label Smoothing (LS), which combines a uniform distribution with the hard ground-truth label, thereby reducing the model’s certainty in the primary class.
+LS has since become prevalent in image recognition [10, 22, 36, 47] and neural machine translation [1, 6], often boosting accuracy and calibration [23].
+Yet subsequent work indicates that LS can overly compress features into tight clusters [15, 28, 41], hindering intra-class variability and transferability [5].
+In parallel, Zhu et al.
+[48] found that LS paradoxically fosters overconfidence in misclassified samples, though the precise mechanism behind this remains uncertain.
 In this paper, we reveal that LS’s training objective inherently contains an error amplification term.
 This term pushes the network to reinforce incorrect predictions with exaggerated certainty, yielding highly confident misclassifications and further compressing feature clusters (Section˜3.1, Table˜1).
-Building on zhu2022rethinking, we characterize “overconfidence” in terms of the model’s top-1 prediction, rather than through conventional calibration metrics.
+Building on Zhu et al.
+[48], we characterize “overconfidence” in terms of the model’s top-1 prediction, rather than through conventional calibration metrics.
 Through our analysis, we further show that punishing the ground-truth logit during misclassification reduces intra-class variation (Table˜2), a phenomenon corroborated by Grad-CAM visualizations (Figure˜2).
@@
 Regularization techniques aim to improve the generalization of deep neural networks by constraining model complexity.
-Classical methods like $\ell_{2}$ [krogh1991simple] and $\ell_{1}$ [zou2005regularization] impose direct penalties on large or sparse weights, while Dropout [Srivastava2014] randomly deactivates neurons to discourage over-adaptation.
-In the realm of loss-based strategies, Label Smoothing (LS) [szegedy2016rethinking] redistributes a fraction of the label probability mass away from the ground-truth class, thereby improving accuracy and calibration [muller2019does].
-Variants such as Online Label Smoothing (OLS) [zhang2021delving] and Zipf Label Smoothing (Zipf-LS) [liang2022efficient] refine LS by dynamically adjusting the smoothed labels based on a model’s evolving predictions.
+Classical methods like $\ell_{2}$ [18] and $\ell_{1}$ [49] impose direct penalties on large or sparse weights, while Dropout [32] randomly deactivates neurons to discourage over-adaptation.
+In the realm of loss-based strategies, Label Smoothing (LS) [34] redistributes a fraction of the label probability mass away from the ground-truth class, thereby improving accuracy and calibration [23].
+Variants such as Online Label Smoothing (OLS) [45] and Zipf Label Smoothing (Zipf-LS) [21] refine LS by dynamically adjusting the smoothed labels based on a model’s evolving predictions.
 However, they do not fully address the fundamental issue that emerges when the ground-truth logit is not the highest one (see Section˜3.1, Table 1).
 Other loss-based regularizers focus on alternative aspects of the predictive distribution.
-Confidence Penalty [pereyra2017regularizing] penalizes the model’s confidence directly, while Logit Penalty [dauphin2021deconstructing] minimizes the global $\ell_{2}$-norm of logits, a technique reported to enhance class separation [kornblith2021better].
+Confidence Penalty [25] penalizes the model’s confidence directly, while Logit Penalty [4] minimizes the global $\ell_{2}$-norm of logits, a technique reported to enhance class separation [15].
 Despite these benefits, Logit Penalty can inadvertently shrink intra-class variation, thereby hampering transfer learning (see Section˜4.1).
@@
 Label Smoothing has also been studied extensively under knowledge distillation.
-For instance, yuan2020revisiting observed that LS can approximate the effect of a teacher–student framework, while shen2021label investigated its role in such pipelines more systematically.
-Additionally, chandrasegaran2022revisiting demonstrated that a low-temperature, LS-trained teacher can notably improve distillation outcomes.
-Concurrently, kornblith2021better showed that LS tightens intra-class clusters in the feature space, diminishing transfer performance.
-From a Neural Collapse perspective [zhou2022all, guo2024cross], LS nudges the model toward rigid feature clusters, as evidenced by the reduced feature variability measured in xu2023quantifying.
+For instance, Yuan et al.
+[43] observed that LS can approximate the effect of a teacher–student framework, while Shen et al.
+[30] investigated its role in such pipelines more systematically.
+Additionally, Chandrasegaran et al.
+[2] demonstrated that a low-temperature, LS-trained teacher can notably improve distillation outcomes.
+Concurrently, Kornblith et al.
+[15] showed that LS tightens intra-class clusters in the feature space, diminishing transfer performance.
+From a Neural Collapse perspective [46, 8], LS nudges the model toward rigid feature clusters, as evidenced by the reduced feature variability measured in Xu and Liu [41].
 Our goal is to overcome LS’s inherent error amplification effect.
@@
     • Inconsistent Regularization: The regularization term lowers $z_{gt}$ yet does not penalize $z_{\textit{max}}$, which further impairs learning.
-These issues with LS on misclassified samples have also been systematically observed in prior work [xia2024understanding].
+These issues with LS on misclassified samples have also been systematically observed in prior work [39].
 By precisely disentangling these two components (regularization vs.
@@
 Ablation Study on LS Components.
-To isolate the effects of each component in LS, we carefully perform a detailed and systematic ablation study on ImageNet-1K using a DeiT-Small model [touvron2021training] without Mixup or CutMix.
+To isolate the effects of each component in LS, we carefully perform a detailed and systematic ablation study on ImageNet-1K using a DeiT-Small model [36] without Mixup or CutMix.
 As indicated in Table˜1, the performance gains from LS stem solely from the regularization term, whereas the error amplification term degrades accuracy.
@@
 For completeness, Appendix B provides the full gradient derivation.
-While [xia2024understanding] conducted a related gradient analysis of the training loss, it focuses specifically on the setting of selective classification, and examines a posthoc logit normalization technique to mitigate confidence calibration issues.
+While [39] conducted a related gradient analysis of the training loss, it focuses specifically on the setting of selective classification, and examines a posthoc logit normalization technique to mitigate confidence calibration issues.
 However, this approach addresses only the overconfidence problem of label smoothing (LS), without tackling representation collapse.
@@
 By contrast, MaxSup uniformly penalizes the top-1 logit, whether the prediction is correct or incorrect, thereby eliminating LS’s error-amplification effect and preserving finer distinctions within each class.
-Table 2 compares intra-class variation ($\bar{d}_{\text{within}}$) and inter-class separability ($R^{2}$[kornblith2021better] for ResNet-50 trained on ImageNet-1K.
+Table 2 compares intra-class variation ($\bar{d}_{\text{within}}$) and inter-class separability ($R^{2}$[15] for ResNet-50 trained on ImageNet-1K.
 Although all investigated regularizers decrease $\bar{d}_{\text{within}}$ relative to a baseline, MaxSup yields the smallest reduction, indicating a stronger retention of subtle within-class diversity—widely associated with enhanced generalization and improved transfer performance.
@@
 4.1.2 Connection to Logit Penalty
-As detailed in Section 3, both Label Smoothing (LS) variants and MaxSup impose penalties directly at the logit level, aligning with the perspective that various regularizers influence a model’s representational capacity via distinct logit constraints [kornblith2021better].
+As detailed in Section 3, both Label Smoothing (LS) variants and MaxSup impose penalties directly at the logit level, aligning with the perspective that various regularizers influence a model’s representational capacity via distinct logit constraints [15].
 Within this family of techniques, Logit Penalty and MaxSup both address the maximum logit, yet diverge fundamentally in their specific methods of regularization.
@@
 4.2.1 Experiment Setup
-Model Training Configurations.We evaluate both convolutional (ResNet[he2016deep], MobileNetV2 [sandler2018mobilenetv2]) and transformer (DeiT-Small [touvron2021training]) architectures on ImageNet [krizhevsky2012imagenet].
+Model Training Configurations.We evaluate both convolutional (ResNet[10], MobileNetV2 [27]) and transformer (DeiT-Small [36]) architectures on ImageNet [17].
 For the ResNet Series, we train for 200 epochs using stochastic gradient descent (SGD) with momentum0.9, weight decay of $1\times 10^{-4}$, and a batch size of 2048.
@@
 We also test ResNet variants on CIFAR-100 with a conventional setup: an initial learning rate of 0.1 (reduced fivefold at epochs 60, 120, and 160), training for 200 epochs with batch size 128 and weight decay $5\times 10^{-4}$.
-For DeiT-Small, we use the official codebase [touvron2021training], training from scratch without knowledge distillation to isolate MaxSup’s contribution.
+For DeiT-Small, we use the official codebase [36], training from scratch without knowledge distillation to isolate MaxSup’s contribution.
 CutMix and Mixup are disabled to ensure the model optimization objective remains unchanged.
-Hyperparameters for Compared Methods.We compare Max Suppression Regularization against a range of LS extensions, including Zipf Label Smoothing[liang2022efficient] and Online Label Smoothing [zhang2021delving].
+Hyperparameters for Compared Methods.We compare Max Suppression Regularization against a range of LS extensions, including Zipf Label Smoothing[21] and Online Label Smoothing [45].
 Where official implementations exist, we adopt them directly; otherwise, we follow the methodological details provided in each respective paper.
@@
 Across all convolutional architectures tested, MaxSup consistently delivers the highest top-1 accuracy among label-smoothing approaches.
-By contrast, OLS [zhang2021delving] and Zipf-LS [liang2022efficient] exhibit less stable gains, suggesting their effectiveness may heavily hinge on specific training protocols.
+By contrast, OLS [45] and Zipf-LS [21] exhibit less stable gains, suggesting their effectiveness may heavily hinge on specific training protocols.
 To reproduce OLS and Zipf-LS, we apply the authors’ official codebases and hyperparameters but do not replicate their complete training recipes (e.g., OLS trains for 250 epochs with a step-scheduled learning rate of 0.1, and Zipf-LS uses 100 epochs with distinct hyperparameters).
@@
 Table 6: Classification on CUB and Cars Datasets.
-Method CUB[cub] Cars[car]
+Method CUB[37] Cars[16]
 Baseline 80.88 90.27
@@
 Fine-Grained Classification.
-Beyond large-scale benchmarks like ImageNet, we further evaluate MaxSup on two fine-grained visual recognition tasks: CUB-200-2011 [cub] and Stanford Cars [car].
+Beyond large-scale benchmarks like ImageNet, we further evaluate MaxSup on two fine-grained visual recognition tasks: CUB-200-2011 [37] and Stanford Cars [16].
 These datasets pose unique challenges due to subtle inter-class differences, which often expose the limitations of standard regularization approaches.
@@
 Long-Tailed Classification.
-To assess the effectiveness of MaxSup under data imbalance, we performed experiments on the CIFAR-10-LT dataset with imbalance ratios of 50 and 100, following the experimental settings described in [tang2020long].
+To assess the effectiveness of MaxSup under data imbalance, we performed experiments on the CIFAR-10-LT dataset with imbalance ratios of 50 and 100, following the experimental settings described in [35].
 The corresponding results are summarized in Table˜7.
@@
 ECE (Corr) 0.145(0.003) 0.158(0.015) 0.260(0.015)
-Corrupted Image Classification To evaluate the effectiveness of MaxSup on out-of-distribution (OOD) settings, we also conducted experiments on CIFAR10-C benchmark [hendrycksbenchmarking] shown in Table˜8 following settings in [heinonen2025robust].
+Corrupted Image Classification To evaluate the effectiveness of MaxSup on out-of-distribution (OOD) settings, we also conducted experiments on CIFAR10-C benchmark [12] shown in Table˜8 following settings in [11].
 Table˜8 reports the performance of MaxSup and Label Smoothing (LS) on this benchmark using ResNet-50 as the backbone.
@@
 We further investigate MaxSup’s applicability to downstream tasks by evaluating its performance on semantic segmentation using the widely adopted MMSegmentation framework.
-[^3] https://github.com/open-mmlab/mmsegmentation Specifically, we adopt the UperNet [xiao2018unified] architecture with a DeiT-Small backbone, trained on ADE20K.
+[^3] https://github.com/open-mmlab/mmsegmentation Specifically, we adopt the UperNet [40] architecture with a DeiT-Small backbone, trained on ADE20K.
 Models pretrained on ImageNet-1K with either MaxSup or Label Smoothing are then fine-tuned under the same cross-entropy objective (Section˜4.2.2).
@@
 By more effectively capturing class boundaries and within-class variability, MaxSup promotes stronger segmentation results, underscoring its potential to deliver features that are both transferable and highly robust.
-Figure 2: Grad-CAM [Selvaraju_2019] visualizations for DeiT-Small models under three training setups: MaxSup (2nd row), Label Smoothing (3rd row), and a baseline (4th row).
+Figure 2: Grad-CAM [29] visualizations for DeiT-Small models under three training setups: MaxSup (2nd row), Label Smoothing (3rd row), and a baseline (4th row).
 The first row shows the original images.
@@
 4.4 Visualization via Class Activation Maps
-To better understand how MaxSup fundamentally differs from Label Smoothing (LS) in guiding model decisions, we employ Gradient-weighted Class Activation Mapping (Grad-CAM) [Selvaraju_2019], which highlights regions most influential for each prediction.
+To better understand how MaxSup fundamentally differs from Label Smoothing (LS) in guiding model decisions, we employ Gradient-weighted Class Activation Mapping (Grad-CAM) [29], which highlights regions most influential for each prediction.
 We evaluate DeiT-Small under three training setups: MaxSup (second row), LS (third row), and a baseline with standard cross-entropy (fourth row).
@@
 Limitations.
-Prior work [muller2019does] notes that LS-trained teachers may degrade knowledge distillation [hinton2015distilling, Hu2021CVPR], and guo2024cross suggests LS accelerates convergence via improved conditioning.
+Prior work [23] notes that LS-trained teachers may degrade knowledge distillation [13, 14], and Guo et al.
+[8] suggests LS accelerates convergence via improved conditioning.
 Examining MaxSup’s potential role in distillation and its overall impact on training dynamics would clarify these underlying effects.
-Recent studies [sukenik2024neural, garrod2024persistence] also show that $\ell_{2}$ regularization biases final-layer features toward low-rank solutions, raising interesting questions about whether MaxSup behaves similarly.
+Recent studies [33, 7] also show that $\ell_{2}$ regularization biases final-layer features toward low-rank solutions, raising interesting questions about whether MaxSup behaves similarly.
 Impact.
@@
 Acknowledgments
-This work was supported by the University of Washington Faculty Startup Fund, the Carwein–Andrews Fellowship, the UW GSFEI Top Scholar Award, and the U.S.
-DOT PacTrans sub-center seed funding program.
-We also thank the anonymous reviewers for their constructive feedback and insightful suggestions, which helped improve the quality and clarity of this paper.
+This work was supported by the University of Washington Faculty Startup Fund, the Carwein–Andrews Fellowship, the UW GSFEI Top Scholar Award, the U.S.
+DOT PacTrans sub-center seed funding program, the DFG Research Unit 5336 - Learning to Sense (L2S), and the ELSA – European Lighthouse on Secure and Safe AI funded by the European Union under grant agreement No.
+101070617.
+Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or European Commission.
+Neither the European Union nor the European Commission can be held responsible for them.
+We thank the anonymous reviewers for their helpful comments.
 NeurIPS Paper Checklist
@@
 Although Label Smoothing and OLS improve ImageNet accuracy, they substantially degrade transfer accuracy compared to MaxSup.
-Following [kornblith2021better], we selected from 45 logarithmically spaced values between $10^{-6}$ and $10^{5}$.
+Following [15], we selected from 45 logarithmically spaced values between $10^{-6}$ and $10^{5}$.
 Notably, the search range is larger than the search range used in Table 3, thus leading to higher overall accuracies on CIFAR10.
@@
 Appendix I Extended Comparison to More Label Smoothing Alternatives
-We have included Confidence Penalty [pereyra2017regularizing] and Adaptive Label Smooothing with Self-Knowledge [lee2022adaptive] for an extended comparison in Table˜13.
+We have included Confidence Penalty [25] and Adaptive Label Smooothing with Self-Knowledge [20] for an extended comparison in Table˜13.
 We follow their recommended hyperparameter settings: weight coefficient is set to 0.1 for confidence penalty, and the checkpoint with the highest validation accuracy checkpoint is treated as teacher for Adaptive Label Smoothing.
-To further address the novelty of our work, we additionally compared our method to recent approaches that also identify and aim to fix the issues of Label Smoothing on misclassified samples [gao2020towards, wei2022mitigatingneuralnetworkoverconfidence].
-Specifically, the method proposed in [wei2022mitigatingneuralnetworkoverconfidence] aims to mitigate overconfidence via logit normalization during training.
+To further address the novelty of our work, we additionally compared our method to recent approaches that also identify and aim to fix the issues of Label Smoothing on misclassified samples [6, 38].
+Specifically, the method proposed in [38] aims to mitigate overconfidence via logit normalization during training.
 With its default hyperparameter settings, it achieves an accuracy of 74.32% on ImageNet with a ResNet-50, which is significantly lower than the 76.91% achieved by standard Label Smoothing.
 This performance aligns with that of Logit Penalty, which similarly minimizes the global $l_{2}$-norm of logits and can struggle to match baseline LS performance.
-We also note that these norm-based methods[gao2020towards, wei2022mitigatingneuralnetworkoverconfidence, dauphin2021deconstructing] are often highly sensitive to hyperparameter choices, which can limit their practical applicability.
+We also note that these norm-based methods[6, 38, 4] are often highly sensitive to hyperparameter choices, which can limit their practical applicability.
 As shown in 13, MaxSup outperforms all these alternatives on ImageNet with ResNet-50.
@@
 Appendix J Visualization of the Learned Feature Space
-To illustrate the differences between Max Suppression Regularization and Label Smoothing, we follow the projection technique of muller2019does.
+To illustrate the differences between Max Suppression Regularization and Label Smoothing, we follow the projection technique of Müller et al.
+[23].
 Specifically, we select three semantically related classes and construct an orthonormal basis for the plane intersecting their class templates in feature space.
@@
 The consistently high accuracy across settings demonstrates the robustness of MaxSup.
-The adaptive $\alpha$ schedule, adopted from [lee2022adaptive], further highlights the method’s integrity and compatibility with principled design choices.
+The adaptive $\alpha$ schedule, adopted from [20], further highlights the method’s integrity and compatibility with principled design choices.
 Table 15: Ablation study on alpha schedules using ResNet50 on ImageNet1K.